### v2

# Computationally Differentially Private Inner-Product Protocols Imply Oblivious Transfer

Abstract
In distributed differential privacy, multiple parties collaborate to analyze their combined data while each party protects the confidentiality of its data from the others. Interestingly, for certain fundamental two-party functions, such as the inner product and Hamming distance, the accuracy of distributed solutions significantly lags behind what can be achieved in the centralized model. For computational differential privacy, however, these limitations can be circumvented using oblivious transfer (used to implement secure multi-party computation). Yet, no results show that oblivious transfer is indeed necessary for accurately estimating a non-Boolean functionality. In particular, for the inner-product functionality, it was previously unknown whether oblivious transfer is necessary even for the best possible constant additive error.
[…]
Keywords: differential privacy; inner product; oblivious transfer
1 Introduction
Differential privacy [DMNS06] is a mathematical privacy guarantee designed to facilitate statistical analyses of databases while ensuring strong guarantees against the leakage of individual-level information. Formally,
[…]
A randomized function (“mechanism”) $\mathsf{M}\colon{\cal X}^{n}\rightarrow{\cal Y}$ is $(\varepsilon,\delta)$-differentially private (in short, $(\varepsilon,\delta)$-${\sf DP}$) if for every $x,x^{\prime}\in{\cal X}^{n}$ that differ on only one entry, and any algorithm $\mathsf{D}\colon{\cal Y}\rightarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}$ (“distinguisher”), it holds that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{D}(\mathsf{M}(x))=1}\right]\leq e^{\varepsilon}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{D}(\mathsf{M}(x^{\prime}))=1}\right]+\delta.$ (1)
When $\delta=0$, we omit it from the notation.
Most research in differential privacy regards the centralized model: users share their data with a central server that applies a differentially private mechanism and releases the results to the (potentially adversarial) world. In many scenarios, however, there is no such trusted server, and the data is distributed among several parties who wish to perform a joint computation while preserving the privacy of each party’s data from the eyes of the others. Constructing multiparty protocols that realize a functionality while preserving ${\sf DP}$ is typically more challenging than creating them in the centralized model; a party’s output should not only be a differentially private given the other parties’ inputs and outputs (as in the centralized model), but also given the other parties’ view in the protocol ([DN04, BNO08]).
A fundamental example that illustrates the hardness of distributed ${\sf DP}$ is the inner-product functionality of two binary datasets: parties $\mathsf{A}$ and $\mathsf{B}$ hold datasets $x\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ and $y\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ (respectively), and wish to estimate the inner product over the integers $\langle x,y\rangle=\sum_{i=1}^{n}x_{i}\cdot y_{i}$ (the correlation between the datasets) using a ${\sf DP}$ protocol. There exists a protocol (specifically, the so-called randomized response) that is $\varepsilon$-${\sf DP}$ and estimates the inner-product up to additive error $O_{\varepsilon}(\sqrt{n})$ ([McGregorMPRTV10, MPRV09]), and this essentially the best error rate we can achieve while maintaining $\varepsilon$-${\sf DP}$: [McGregorMPRTV10, HaitnerMST22] showed that any $\varepsilon$-${\sf DP}$ protocol for the inner-product, even when allowing any $\delta\in o(1/n)$, must have an additive error of $\Omega_{\varepsilon}(\sqrt{n})$. This is in contrast to the centralized model, where the simple $\varepsilon$-${\sf DP}$ Laplace mechanism $\mathsf{M}(x,y)=\langle x,y\rangle+{\rm Lap}(2/\varepsilon)$ has only additive error $O_{\varepsilon}(1)$.
To achieve better accuracy, [BNO08, MPRV09] considered protocols that only guarantee computational differential privacy (${\sf CDP}$): a computational analog of ${\sf DP}$ where Section 1 is relaxed so that Equation 1 only holds for computationally efficient (ppt) algorithms $\mathsf{D}$, and not necessarily for any algorithm $\mathsf{D}$. This relaxation is very powerful. For example, if we assume the existence of the cryptographic primitive oblivious transfer (in short, ${\sf OT}$, [Rabin81]), then in any distributed setting (even for more than two parties), we can achieve the accuracy of a centralize ${\sf DP}$ mechanism by emulating it using secure multiparty computation ([BNO08, DKMMN06]). In particular, using ${\sf OT}$, two parties can securely emulate the (centralized) Laplace mechanism to achieve an $\varepsilon$-${\sf CDP}$ protocol with additive error $O_{\varepsilon}(1)$. This leads to the following fundamental question:
[…]
Can we construct a two-party ${\sf CDP}$ protocol that estimates the inner product with a small additive error using a cryptographic primitive that is weaker than ${\sf OT}$?
[…]
*HaitnerMST22 made partial progress on this matter, showing that any $(\varepsilon,\frac{1}{n^{2}})$-${\sf CDP}$ protocol with non-trivial accuracy guarantee of $o_{\varepsilon}(\sqrt{n})$ must use a cryptographic primitive that is at least as strong as key agreement (in short, ${\sf KA}$). However, although ${\sf KA}$ and ${\sf OT}$ are both public-key cryptography primitives, ${\sf KA}$ is believed to be a weaker primitive ([GertnerKMRV00]) and does not suffice for secure MPC. Even for ${\sf CDP}$ protocols that estimate the inner-product with very high accuracy of $O_{\varepsilon}(1)$, it was previously unknown whether ${\sf OT}$ is indeed necessary.
1.1 Our Result
We prove the first equivalence between ${\sf OT}$ and ${\sf CDP}$ two-party protocols that estimate the inner-product with small additive error, and in particular, additive error up to $O_{\varepsilon}(n^{1/6})$.
Theorem 1.1 (Main result, informal).
An $\mathopen{}\mathclose{{}\left(\varepsilon,\frac{1}{3n}}\right)$-${\sf CDP}$ two-party protocol that estimates the inner product over $\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ up to an additive error $n^{1/6}/(c\cdot e^{c\varepsilon})$ with probability $0.999$ (for some universal constant $c>0$), can be used to construct an ${\sf OT}$ protocol. [^1] Theorem 1.1 is proven via a uniform and black-box reduction. Thus, although the security of ${\sf CDP}$ and ${\sf OT}$ is commonly defined with respect to non-uniform adversaries, Theorem 1.1 also implies a reduction from ${\sf OT}$ that is secure against uniform adversaries to such mildly accurate inner-product protocols that are ${\sf CDP}$ with respect to uniform distinguishers.
In particular, Theorem 1.1 implies that without assuming the existence of ${\sf OT}$, it is impossible to construct a ${\sf CDP}$ protocol for the inner-product with a constant, or even sub-polynomial, additive error. It follows that a protocol with a small polynomial additive error is equivalent to a protocol with a constant error. That is because using such a mildly accurate protocol, we can construct by Theorem 1.1 an oblivious transfer, and then use it to perform a secure two-party computation of the Laplace mechanism.
[…]
There exists a universal constant $c>0$ such that the following holds: assume that an $\mathopen{}\mathclose{{}\left(\varepsilon,\frac{1}{3n}}\right)$-${\sf CDP}$ two-party protocol that estimates the inner product over $\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ up to an additive error $n^{1/6}/(c\cdot e^{c\varepsilon})$ with probability $0.999$ exists, then there exists an $\mathopen{}\mathclose{{}\left(\varepsilon,\operatorname{neg}(n)}\right)$-${\sf CDP}$ protocol that estimates the inner product up to an additive error $20/\varepsilon$ with probability $0.999$.
Theorem 1.1 is the first equivalence result between ${\sf OT}$ to any ${\sf CDP}$ protocol that estimates a natural, non-Boolean functionality. Prior works only provide such equivalences for ${\sf CDP}$ protocols that estimate the (Boolean) XOR functionality ([GKMPS16, HMSS19]), a less interesting functionality in the context of ${\sf DP}$, as even in the centralized model, any ${\sf DP}$ algorithm for estimating XOR must incur an error close to one-half. In contrast, the inner product has a significantly larger gap between the accuracy achievable in two-party ${\sf DP}$ and in ${\sf CDP}$. A proof overview of Theorem 1.1 is given in Section 2.
1.1.0.1 Comparison with [HaitnerMST22].
We remark that our result is not strictly stronger than [HaitnerMST22]. First, [HaitnerMST22] covered the whole non-trivial accuracy regime up to $O_{\varepsilon}(\sqrt{n})$ while our result is only limited to $O_{\varepsilon}(n^{1/6})$. Second, [HaitnerMST22]’s result holds even for low confidence regimes, while Theorem 1.1 requires high accuracy confidence of $0.999$. But in the natural high-accuracy high-confidence regimes, our result has several advantages beyond the necessity of a stronger primitive. First, [HaitnerMST22]’s reduction only works for ${\sf CDP}$ two-party protocols that provide the inner-product estimation within the transcript, while our reduction in Theorem 1.1 just assumes that (at least) one of the parties sees the estimation. Second, and perhaps surprisingly, the proof of Theorem 1.1 is conceptually simpler, even though it provides reduction from a stronger primitive. We provide more details in Section 2 (see 2.0.0.3).
[…]
Very roughly, in the secrecy analysis of [HaitnerMST22]’s ${\sf KA}$ protocol, they had to prove (in a constructive way) that the inner product of a strong Santha Vazirani (SV) source [McGregorMPRTV10] with a uniformly random seed is a good condenser, even when the seed and the source are dependent. When the seed and the source are independent, [McGregorMPRTV10] already proved that the inner product is not only a good condenser for strong SV sources, but also a good extractor for (standard) SV sources [SV87], which apart of their non-constructive technique and the independency assumption, is a stronger result. However, [HaitnerMST22]’s dependency between the source and the seed resulted in a significant technical challenge, which made the proof conceptually harder. In contrast, in our ${\sf OT}$ security analysis, since we only provide secrecy guarantees from the point of views of the parties (and not from the point of view of an external observer that only sees the transcript), we bypass this dependency issue. See more details in Section 2.
1.2 Open Questions
In this paper, we show that a $\mathopen{}\mathclose{{}\left(\varepsilon,\frac{1}{3n}}\right)$-${\sf CDP}$ protocol that estimates the inner product over $\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ up to an additive error $O_{\varepsilon}(n^{1/6})$ is equivalent to oblivious transfer (${\sf OT}$). For the inner-product functionality, it is still left open to close the gap up to any non-trivial accuracy of $O_{\varepsilon}(\sqrt{n})$. [HaitnerMST22] closed this gap w.r.t. a weaker notion of ${\sf CDP}$ against external observer [^2] A protocol is ${\sf CDP}$ against external observer, if the transcript of the execution is a ${\sf CDP}$ function of the input datasets, but not necessarily the views of the parties as in the standard notion. by showing that key agreement (${\sf KA}$) is necessary and sufficient for any non-trivial accuracy guarantee, which in particular implies that ${\sf KA}$ is also necessary for the standard notion of ${\sf CDP}$. But since ${\sf KA}$ is strictly weaker than ${\sf OT}$ (under block-box reductions), we still do not know what is the right answer for additive errors in the range between $n^{1/6}$ and $\sqrt{n}$.
Beyond characterizing the inner-product functionality, the main challenge is to extend this understanding to other ${\sf CDP}$ distributed computations. For some functionalities, e.g., Hamming distance, we have a simple reduction to the inner-product functionality. But finding a more general characterization that captures more (or even all) functionalities, remains open.
1.3 Paper Organization
In Section 2, we give a high-level proof of Theorem 1.1, and in Section 3, we discuss additional related works. Notations, definitions, and general statements used throughout the paper are given in Section 4. Our main formal theorems are stated in Section 5, and the oblivious transfer protocol together with its security proof are given in Section 6. In Section 3 we discuss additional related works. The proof of our main result relies on a technical tool that is proven in Appendix B. Other missing proofs appear in Appendix C.
2 Proof Overview
In this section, we give an overview of our proof techniques. Our starting point is the key-agreement protocol of [HaitnerMST22]. As mentioned above, [HaitnerMST22] showed that a ${\sf CDP}$ protocol $\Pi$ for computing the inner product functionality implies the existence of key-agreement (${\sf KA}$). To prove that, they used $\Pi$ to construct a (weak) ${\sf KA}$ protocol, in which the parties $\mathsf{A}$ and $\mathsf{B}$ interact as follows:
  1. 1.
    In the first step, $\mathsf{A}$ and $\mathsf{B}$ choose random inputs $x\in\mathopen{}\mathclose{{}\left\{-1,1}\right\}^{n}$ and $y\in\mathopen{}\mathclose{{}\left\{-1,1}\right\}^{n}$, respectively.
  2. 2.
    The parties interact using $\Pi$ to get approximation $z$ of $\langle x,y\rangle$.
  3. 3.
    Next, $\mathsf{A}$ chooses a random string $r\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}$, and sends $r,x_{r}=\mathopen{}\mathclose{{}\left\{x_{i}\colon r_{i}=1}\right\}$ to $\mathsf{B}$. $\mathsf{B}$ replies with $y_{-r}=\mathopen{}\mathclose{{}\left\{y_{i}\colon r_{i}=0}\right\}$.
  4. 4.
    Finally, $\mathsf{A}$ computes and outputs ${\rm out}_{\mathsf{A}}=\langle x_{-r},y_{-r}\rangle$, and $\mathsf{B}$ outputs ${\rm out}_{\mathsf{B}}=z-\langle x_{r},y_{r}\rangle$.
The first observation made by [HaitnerMST22] is that, by linearity of the inner product, $\langle x_{-r},y_{-r}\rangle+\langle x_{r},y_{r}\rangle=\langle x,y\rangle$, and thus $\mathopen{}\mathclose{{}\left|{\rm out}_{A}-{\rm out}_{B}}\right|=\mathopen{}\mathclose{{}\left|z-\langle x,y\rangle}\right|$, which is small assuming $\Pi$ is accurate. Moreover, it was shown that this protocol can be amplified into a full-fledged ${\sf KA}$ protocol if any efficient adversary $\mathsf{E}$ cannot approximate ${\rm out}_{A}$ within a small additive distance, given the transcript of the weak ${\sf KA}$ protocol. The main technical part in [HaitnerMST22] is to show that this is indeed the case in the above protocol.
In this work, we are interested in constructing a stronger primitive, namely oblivious transfer (${\sf OT}$). In contrast to a ${\sf KA}$ protocol in which security must hold against external observers that only see the transcript, in ${\sf OT}$, the security needs to hold against an adversary that has access to the full view of one of the parties. Our first observation is that the original ${\sf KA}$ protocol of [HaitnerMST22], as is, cannot be used to construct any form of even “weak” oblivious transfer. In more details, it is possible to show that for some carefully chosen ${\sf CDP}$ and accurate protocol $\Pi$, the joint view of the parties in the above protocol can be simulated using a trivial protocol, without using $\Pi$ at all. [^3] Here a protocol $\Pi^{\prime}$ simulates $\Pi$ if $\Pi^{\prime}$ generates the joint view of the parties $\mathsf{A}$, $\mathsf{B}$ in $\Pi$, without revealing any other information to the parties. We say that the simulation $\Pi^{\prime}$ is trivial if it does not rely on any nontrivial protocol or cryptographic assumptions. Importantly, the simulation may reveal new information (that is already known for the internal parties $\mathsf{A}$ and $\mathsf{B}$) to an external observer, and thus it doesn’t imply that the protocol is not a ${\sf KA}$ protocol. See Appendix A for a concrete example. Since there is no ${\sf OT}$ without computational assumptions, this leads to the conclusion that the ${\sf KA}$ protocol of [HaitnerMST22] could not be used as a subroutine (in a black-box manner) to construct ${\sf OT}$. Looking ahead, our solution is to inject carefully placed noise to the interaction, in a manner that the parties cannot simulate without revealing private information. As we will see shortly, this enables us to bypass the limitations of the original construction of [HaitnerMST22] and to achieve ${\sf OT}$.
2.0.0.1 Weak Erasure Channel.
Before describing our construction, let us explain the properties that we need the construction to fulfill. Similarly to [HaitnerMST22], we will start with constructing a weak ${\sf OT}$ protocol, and then we will show how to amplify it into a full-fledged ${\sf OT}$ protocol. By the work of [Wullschleger09], to achieve OT, it is sufficient to construct a weak erasure channel (WEC). [^4] A weak erasure channel is a weak version of Rabin’s ${\sf OT}$ [Rabin81], which is known to be equivalent to (${2\choose 1}$-${\sf OT}$. In this work, we construct a protocol with a weaker security guarantee, which we call approximate weak erasure channel (AWEC), and show that such a protocol also suffices for constructing ${\sf OT}$. Informally, an $(\ell,\alpha,p,q)$-AWEC is a no-input protocol between $\mathsf{A}$ and $\mathsf{B}$, party $\mathsf{A}$ outputs a number $O_{\mathsf{A}}\in[-n,n]$, and party $\mathsf{B}$ either outputs a number $O_{\mathsf{B}}\in[-n,n]$, or an erasure symbol $O_{\mathsf{B}}=\bot$. We additionally require that:
  • Erasure happens with probability $1/2$. Namely, ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[O_{\mathsf{B}}=\perp}\right]=1/2$.
  • When there is no erasure, $\mathsf{A}$ and $\mathsf{B}$ approximately agree. More formally,
    ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|O_{A}-O_{B}}\right|>\ell\mid O_{B}\neq\bot}\right]\leq\alpha.$
  • When an erasure occurs, $\mathsf{B}$ cannot predict the value of $O_{A}$: let $V_{B}$ be the view of $\mathsf{B}$ in the protocol, then for any efficient $\widehat{\mathsf{B}}$
    ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|\widehat{\mathsf{B}}(V_{B})-O_{A}}\right|\leq 1000\ell\mid O_{B}=\bot}\right]\leq q.$
    (i.e., if $O_{B}=\bot$, Bob can’t estimate the value of $O_{A}$ with error $\leq 1000\ell$).
  • Lastly, we require that $\mathsf{A}$ cannot tell when an erasure occurs: let $V_{A}$ be the view of $\mathsf{A}$ in the protocol, then for any efficient $\widehat{\mathsf{A}}$
    $\mathopen{}\mathclose{{}\left|{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\widehat{\mathsf{A}}(V_{A})=1\mid O_{B}\neq\perp}\right]-{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\widehat{\mathsf{A}}(V_{A})=1\mid O_{B}=\perp}\right]}\right|\leq p.$
See Section 4.4.4 in Section 4.4.4 for the formal definition of AWEC. We show that when $p+q+\alpha\ll 1$, an AWEC protocol can be amplified into an ${\sf OT}$. We do this by showing a reduction from AWEC to WEC and then applying the amplification result of [Wullschleger09] to get a full-fledged ${\sf OT}$.
2.0.0.2 The Construction.
We next describe how to use a ${\sf CDP}$ protocol $\Pi$ computing the inner-product to construct an AWEC. Our accuracy requirement from $\Pi$ is that on random inputs, $\Pi$ output is within distance $\ell\ll n^{1/6}$ from the inner-product with probability at least $0.999$. That is,
${\mathrm{Pr}}_{x,y\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n},z\leftarrow\Pi(x,y)}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|z-\langle x,y\rangle}\right|\leq\ell}\right]\geq 0.999.$
The construction is similar to the construction of [HaitnerMST22], with the key difference that with probability $1/2$, $\mathsf{B}$ sends a noisy version of its vector to $\mathsf{A}$. Informally, this random noise “erases” the information $\mathsf{B}$ has on $\mathsf{A}$’s output.
[…]
Operation: 1. $\mathsf{A}$ samples $x\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$, and $\mathsf{B}$ samples $y\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$. 2. $\mathsf{A}$ and $\mathsf{B}$ interact according to $\Pi$ using inputs $x$ and $y$ resp., to get output $z\approx\langle x,y\rangle\pm\ell$. 3. $\mathsf{A}$ samples $r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}$ and sends $(r,x_{r}=\mathopen{}\mathclose{{}\left\{x_{i}\colon r_{i}=1}\right\})$ to $\mathsf{B}$. 4. $\mathsf{B}$ samples a random bit $b\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}$ and acts as follows: (a) If $b=0$ (“non-erasure”), it sends $y_{-r}=\mathopen{}\mathclose{{}\left\{y_{i}\colon r_{i}=0}\right\}$ to $\mathsf{A}$ and outputs $o_{B}=z-\langle x_{r},y_{r}\rangle$. (b) Otherwise ($b=1$, “erasure”), it performs the following steps: Samples $k$ uniformly random indices $i_{1},\ldots,i_{k}\leftarrow[n]$, where $n^{1/3}\gg k\gg e^{\varepsilon}\cdot\ell^{2}$. Compute $\widetilde{y}=(\widetilde{y}_{1},\ldots,\widetilde{y}_{n})$ where $\widetilde{y}_{i}\leftarrow\begin{cases}{\cal U}_{\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}}&i\in\mathopen{}\mathclose{{}\left\{i_{1},\ldots,i_{k}}\right\}\\ y_{i}&\text{otherwise}\end{cases}$, Send $\widetilde{y}_{-r}=\mathopen{}\mathclose{{}\left\{\widetilde{y}_{i}\colon r_{i}=0}\right\}$ to $\mathsf{A}$, and Output $o_{B}=\perp$. 5. Denote by $\widehat{y}_{-r}$ the message $\mathsf{A}$ received from $\mathsf{B}$. Then $\mathsf{A}$ outputs $o_{A}=\langle x_{-r},\widehat{y}_{-r}\rangle$.
That is, with probability $1/2$, $\mathsf{B}$ decides to make an erasure by changing the value of $k$ random bits of $y$ before sending $y_{-r}$ to $\mathsf{A}$. Let $\mu=\widetilde{y}-y$ be the noise vector. Then importantly, the crux of the above protocol is that the noise $\langle\mu_{-r},x_{-r}\rangle=\langle\widetilde{y}_{-r},x_{-r}\rangle-\langle y_{-r},x_{-r}\rangle$ cannot be simulated by $\mathsf{A}$ or $\mathsf{B}$ without revealing more information on the input of the other party. We next give an overview of the analysis of the above construction.
2.0.0.3 Analysis.
We define the following random variables with respect to a random execution of 2.0.0.2. Let $X,Y$ be inputs of $\mathsf{A}$ and $\mathsf{B}$ to $\Pi$, respectively, and let $V_{A}$ and $V_{B}$ be their view in the interaction according to $\Pi$. Let $R,I_{1},\dots,I_{k},\widetilde{Y}$ and $\widehat{Y}$ be the random variables taking the values of $r,i_{1},\dots,i_{k},\widetilde{y}$ and $\widehat{y}$, respectively. Finally, let $\widehat{V}_{A}=(V_{A},R,\widehat{Y}_{-R})$ and $\widehat{V}_{B}=(V_{B},R,X_{R},I_{1},\dots,I_{k},\widetilde{Y},\widehat{Y})$ be the parties’ views in 2.0.0.2, and $O_{A},O_{B}$ their outputs, respectively.
It is evident by the definition of the protocol that erasure indeed occurs with probability $1/2$. Moreover, similarly to the correctness of the key-agreement protocol, and by the accuracy of $\Pi$, when there is no erasure, $O_{A}$ and $O_{B}$ are at most $\ell$ apart with probability at least $0.999$. Thus, it remains to prove that $\mathsf{B}$ cannot approximate $O_{A}$ when erasure occurs, and that $\mathsf{A}$ cannot learn whether it happened. To simplify the analysis in this proof overview, in the following, we assume that $\Pi$ is a $(\varepsilon,0)$-CDP protocol, [^5] While $(\varepsilon,0)$-CDP (that is, setting $\delta=0$) doesn’t make sense in the computational setting, for the sake of simplicity, we ignore this subtly in this informal overview. for a small constant $\varepsilon>0$.
Note that while $\mathsf{B}$ learns the same information on $\mathsf{A}$’s input in the protocol in both the erasure and non-erasure cases, $\mathsf{A}$’s output $O_{A}$ is different in each case; when there is no erasure, $O_{A}$ is predictable (up to an additive error $\ell$) given $\mathsf{B}$’s view $V_{B}$. When there is an erasure, we claim that the privacy guarantee of $\Pi$ implies that $O_{A}$, given $V_{B}$, is unpredictable up to an additive error $k^{1/2}\gg\ell$ (more details in Section 2.1). The “magic” of the protocol is in the fact that, for $\ell\ll n^{1/6}$, $\mathsf{B}$ can control $\mathsf{A}$’s output, and manipulate it to be unpredictable (from $\mathsf{B}$’s point of view), without $\mathsf{A}$ noticing the difference. Proving the latter property (hidden erasure) is the main technical challenge of this work, and we provide an overview of the proof in Section 2.2. By combining the two requirements on $k$, we obtain that our protocol is indeed ${\sf AWEC}$.
[…]
Before discussing the security proof of our protocol, we would like to highlight an essential difference between the analysis of our protocol and [HaitnerMST22]’s ${\sf KA}$ protocol. While the primitive we construct here, ${\sf OT}$, is a stronger primitive than ${\sf KA}$, certain aspects of the proof become more manageable in our analysis. This is because in ${\sf OT}$, we need to consider an internal attacker, while in ${\sf KA}$ the security needs to hold against an external observer. In more detail, [HaitnerMST22] requires showing that security holds against an adversary who sees $R,X_{R}$, and $Y_{-R}$. This form of leakage creates a dependency between the seed $R$ and both the leakage on $X$, $X_{R}$, and the leakage of $Y$, $Y_{-R}$, which makes the analysis challenging. In contrast, in our setting, the attacker sees one of the vectors, $X$ or $Y$, entirely. For example, $\mathsf{A}$ gets to see $X,R$ and $Y_{-R}$, and therefore there is no dependency between the leakage on $X$ and $R$. This fact helps us to bypass this dependency issue and makes this part of our analysis more manageable.
2.1 Erasures do Erase
To see why “erasures” in 2.0.0.2 do erase the value of $O_{A}$ from $\mathsf{B}$’s point of view, let’s first consider a different version of 2.0.0.2, where in the erasure case, party $\mathsf{B}$, instead of changing $k$ random bits in $Y_{-R}$, changes all the bits to random ones. (Namely, in the erasure case, $\widehat{Y}_{-R}$ is a uniformly random string, independent of $X$). Now, the question is, can $\mathsf{B}$ now predict $O_{A}=\langle X_{-R},\widehat{Y}_{-R}\rangle$? From $\mathsf{B}$’s point of view, the privacy guarantee of the protocol $\Pi$ implies that $X_{-R}$ must be “unpredictable”, and in [McGregorMPRTV10, HaitnerMST22], it is called “strong Santha Vazirani (SV)” source. [McGregorMPRTV10] showed that the inner product of an $n$-size strong SV source with a random string, even if the string is known, is unpredictable up to an additive error of $\approx\sqrt{n}$, and [HaitnerMST22] extended this result to the computational case. Therefore, if $\mathsf{B}$ decides to erase in this version, it cannot estimate $O_{A}=\langle X_{-R},\widehat{Y}_{-R}\rangle$ with an error smaller than $\approx\sqrt{n}$, which is much larger than $\ell$, its estimation of $\langle X_{-R},\widehat{Y}_{-R}\rangle$ in the non-erasure case. The problem with this version is that $\mathsf{A}$ can distinguish between the erasure and non-erasure cases. Hence, in the actual protocol, the number of bits that $\mathsf{B}$ changes in the erasure case is set to some smaller parameter $k$. The point is that the same argument we applied to the previous version still holds; if $\mathsf{B}$ changes $k$ random values in $Y_{-R}$ (call this $k$-size set of indices by ${\cal H}$), then $O_{A}$ would be a linear function of $\langle X_{{\cal H}},\widehat{Y}_{{\cal H}}\rangle$, and since $\mathsf{B}$ cannot estimated $\langle X_{{\cal H}},\widehat{Y}_{{\cal H}}\rangle$ with error smaller than $\approx\sqrt{k}$ (from the same arguments above), it cannot do it for $O_{A}$ as well.
In more detail, assume towards a contradiction that there exists an efficient $\widehat{\mathsf{B}}$, that given $\widehat{V}_{B}$ and conditioned on $O_{B}=\bot$ (equivalently, on $b=1$) can approximate $O_{A}$ within a small distance. Namely, assume that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|\widehat{\mathsf{B}}(\widehat{V}_{B})-O_{A}}\right|\leq 1000\ell\mid O_{B}=\bot}\right]>q$
for some small constant $q$. Our goal is to construct, using $\widehat{\mathsf{B}}$, an algorithm $\widetilde{\mathsf{B}}$ that breaks the CDP guarantee of $\Pi$.
Building on [McGregorMPRTV10, HaitnerMST22], to break the CDP property of the protocol $\Pi$, it is enough to construct an algorithm that approximates the inner product of (a subset of) $X$ with a random vector $S$. That is, we want to construct an algorithm $\widetilde{\mathsf{B}}$ that for a random subset ${\cal H}\subseteq[n]$ of size $k/4$ and for a random vector $S\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{k/4}$, approximates the inner product $\langle X_{{\cal H}},S\rangle$ within small additive distance. More formally, we want an algorithm $\widetilde{\mathsf{B}}$ such that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|\widetilde{\mathsf{B}}(V_{B},{\cal H},S,X_{-{\cal H}})-\langle X_{{\cal H}},S\rangle}\right|\leq\sqrt{\mathopen{}\mathclose{{}\left|{\cal H}}\right|/e^{\varepsilon}}}\right]>\alpha.$
The work of [McGregorMPRTV10] implies that in the information-theoretic settings, such an algorithm contradicts the privacy guarantee of $\Pi$666[McGregorMPRTV10] showed that when ${\cal H}=[n]$, approximating the inner product contradicts DP. Extending this result for any subset ${\cal H}$ is not hard. [HaitnerMST22] extended this result to the computational settings, by giving a constructive proof. Thus, such an algorithm $\widetilde{\mathsf{B}}$ is enough to break the assumed CDP property of $\Pi$.
[…]
To see why, let ¯=$\bar{{\cal H}}=\mathopen{}\mathclose{{}\left\{i\colon R_{i}=0}\right\}\setminus{\cal H}$, and let $T=\widehat{\mathsf{B}}(\widehat{V}_{B})-\langle X_{\bar{{\cal H}}},\widetilde{Y}_{\bar{{\cal H}}}\rangle$. We claim that $T$ is a good approximation of $\langle X_{{\cal H}},S\rangle$. Indeed,
$T-\langle X_{{\cal H}},S\rangle$ $=T-(\langle X_{-R},\widetilde{Y}_{-R}\rangle-\langle X_{\bar{{\cal H}}},\widetilde{Y}_{\bar{{\cal H}}}\rangle)$
$=(T+\langle X_{\bar{{\cal H}}},\widetilde{Y}_{\bar{{\cal H}}}\rangle)-\langle X_{-R},\widetilde{Y}_{-R}\rangle$
$=\widehat{\mathsf{B}}(V_{B})-\langle X_{-R},\widetilde{Y}_{-R}\rangle.$
Therefore, $\mathopen{}\mathclose{{}\left|T-\langle X_{{\cal H}},S\rangle}\right|\leq\sqrt{\mathopen{}\mathclose{{}\left|{\cal H}}\right|/e^{\varepsilon}}$ if $\mathopen{}\mathclose{{}\left|\widehat{\mathsf{B}}(V_{B})-\langle X_{-R},\widetilde{Y}_{-R}\rangle}\right|\leq 1000\ell$ (note that by our choice of $k$, $\ell\ll\sqrt{\mathopen{}\mathclose{{}\left|{\cal H}}\right|/e^{\varepsilon}}=\sqrt{k/4e^{\varepsilon}}$).
Thus, we only need to show an algorithm $\widetilde{\mathsf{B}}$ that, given $V_{B},{\cal H},S,X_{-{\cal H}}$ as input, evaluates the value of $T$. The main issue left is that while $\widetilde{\mathsf{B}}$ only gets $V_{B}$, and to evaluate $T$ we need to execute $\widehat{B}$ that needs to get an entire view $\widehat{V}_{B}$ of $\mathsf{B}$ in 2.0.0.2. However, this can be easily simulated. Indeed, $\widetilde{\mathsf{B}}$ can simply sample $R,I_{1},\dots,I_{k}$ and $\widetilde{Y}$ condition on the values of ${\cal H},S$ (and $O_{B}=\bot$), to get the exact distribution of $\widehat{V}_{B}$.
2.2 Erasures are Hidden
This part of the proof is more involved and is our main new technical part. The goal is to show that $\widehat{\mathsf{A}}$ cannot tell when an erasure happens; equivalently, that $\widehat{\mathsf{A}}$ cannot distinguish $Y_{-R}$ from the noisy version $\widetilde{Y}_{-R}$ with too good probability. For this end, we show that an algorithm $\widehat{\mathsf{A}}$ that (given $V_{A}$) can distinguish $Y_{-R}$ from the noisy version $\widetilde{Y}_{-R}$ with a small but constant probability, can be used to break ${\sf CDP}$.
[…]
The first step is to notice that, by a rather standard hybrid argument, it is enough to show that $\widehat{\mathsf{A}}$ cannot distinguish with advantage $\Theta(1/k)$ between $Y$ and a vector $Y^{\prime}$ which is identical in all but one random coordinate. That is, let $I\leftarrow[n]$ be a random index, and let $Y^{(I)}=(Y_{1},\dots,Y_{I-1},-Y_{I},Y_{I+1},...,Y_{n})$ be the vector that is identical to $Y$ in all entries except the $I$-th one. Then it is enough to show that if
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\widehat{\mathsf{A}}(V_{A},R,Y_{-R})=1}\right]-{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\widehat{\mathsf{A}}(V_{A},R,Y^{(I)}_{-R})=1}\right]\geq 1/k,$
then $\widehat{\mathsf{A}}$ can be used to break the $\varepsilon$-${\sf CDP}$ of $\Pi$.
We highlight that the key difficulty in proving the above lies in the fact that the distinguishing advantages are very small. Indeed, an algorithm $\widehat{A}$ that distinguishes $Y_{-R}$ from $Y^{(I)}_{-R}$ with a large enough constant advantage (proportional to $\varepsilon$) already contradicts the privacy requirement of $\Pi$. However, in our setup, the distinguishing advantage is much smaller than $\varepsilon$, which makes the task of contradicting $\varepsilon$-CDP challenging. We next emphasize two crucial points that make $\widehat{\mathsf{A}}$ useful (and our design for our AWEC protocol, 2.0.0.2, was guided by these two observations).
The importance of a random index
We note that it is crucial that the index $I$ is chosen at random and that it is unknown to $\widehat{\mathsf{A}}$. To see that, let $\mathsf{M}$ be a DP mechanism and let $Y\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ be an input. We claim that it can be the case that an adversary can distinguish between $Y$ and $Y^{(i)}$ given $\mathsf{M}(Y)$ with advantage $\approx\varepsilon$, for any fixed index $i$. Indeed, it is possible that $\mathsf{M}(Y)$ outputs a noisy estimation of $Y_{i}$, which is correct with probability $\approx 1/2+\varepsilon$. Such an estimation will not contradict differential privacy, but will be useful to distinguish between $Y$ and $Y^{(i)}$. We resolve this issue by adding the noise in a random position. To prove that the existence of such noise (in a random position) is indeed hidden, we developed and proved Section 2.2 which is our main technical lemma.
The importance of a random subset
Next, we claim that it is crucial that $\widehat{\mathsf{A}}$ only gets $Y_{-R}$, a random subset of $Y$. Indeed, as in the previous example, the mechanism $\mathsf{M}(Y)$ can output a noisy estimation of the parity of the number of ones in $Y$. As in the previous example, such an estimation does not contradict DP, but is enough to distinguish $Y$ and $Y^{(I)}$ by considering the number of ones in the input compared to the estimation. We deal with such a global information by revealing to $\widehat{\mathsf{A}}$ only a subset of $Y$. This makes the global information not useful.
Thus, in the proof, we need to leverage the fact that $\widehat{\mathsf{A}}$ distinguishes $Y_{-R}$ from $Y^{(I)}_{-R}$ for a random index $I$ and a random subset $R$. We use such an algorithm $\widehat{\mathsf{A}}$ to construct an algorithm $\mathsf{G}$ that given $Y_{-I}=(Y_{1},\ldots,Y_{I-1},Y_{I+1},\ldots,Y_{n})$, predicts the value of $Y_{I}$. In more detail, we prove the following lemma, which is the technical core of our proof.
[…]
Let $(Z,W)$ be jointly distributed random variables, let $R\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}$ and $I\leftarrow[n]$. Let $\mathsf{A}$ be an efficient algorithm that satisfies
$\mathopen{}\mathclose{{}\left|{\mathrm{E}}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z_{R},W)-\mathsf{A}(R,Z^{(I)}_{R},W)}\right]}\right|\geq 1/k,$
for $Z^{(I)}=(Z_{1},\ldots,Z_{I-1},-Z_{I},Z_{I+1},\ldots,Z_{n})$. Then, there exists an efficient algorithm $\mathsf{G}$ that outputs a value in $\mathopen{}\mathclose{{}\left\{1,-1,\bot}\right\}$, such that the following holds.
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}(I,Z_{-I},W)=-Z_{I}}\right]\leq\frac{k^{2}}{n},\ and,\ {\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}(I,Z_{-I},W)=Z_{I}}\right]\geq 1/k-k^{2}/n.$
Back to our setting, by taking $\mathsf{A}(R,Z_{R},W)=\widehat{\mathsf{A}}((1,\ldots,1)-R,Z_{R},W)$, $Z=Y$ and $W=V_{A}$ in Section 2.2, we get an algorithm $\mathsf{G}$ such that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}(I,Y_{-I},V_{A})=-Y_{I}}\right]\leq\frac{k^{2}}{n},\ and,\ {\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}(I,Y_{-I},V_{A})=Y_{I}}\right]\geq 1/k-k^{2}/n.$
We highlight that the gap between the probability of $\mathsf{G}$ to output $Y_{I}$ to its probability to output $-Y_{I}$ is smaller than $1/k$, and that it is not enough to break DP alone. The point is that we have a bound on the probability of $\mathsf{G}$ to output $-Y_{I}$, and that in DP, we measure the difference in probabilities using multiplicative distance. Thus, when
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}(I,Y_{-I},V_{A})=Y_{I}}\right]>e^{\varepsilon}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}(I,Y_{-I},V_{A})=-Y_{I}}\right]$ (2)
the algorithm $\mathsf{G}$ does imply contradiction to DP. To make sure that Equation 2 holds, we need to have
$1/k-k^{2}/n>e^{\varepsilon}\cdot k^{2}/n,$
which is the reason we need to take $k\ll n^{1/3}$. In the rest of this section, we give a high-level proof sketch of Section 2.2.
2.3 Proof Overview for Section 2.2
In this part, we explain the ideas behind the proof of Section 2.2. We start by describing the algorithm $\mathsf{G}$.
2.3.0.1 The predictor $\mathsf{G}$.
The construction of $\mathsf{G}$ leverages the fact that $\mathsf{A}$ succeed on a random index $I$ while only seeing a subset of $Y$. Recall that on input $(I,Z_{-I},W)$, the goal of $\mathsf{G}$ is to predict the value of $Z_{I}$. Towards this end, let $Z^{b}=(Z_{1},\dots,Z_{I-1},b,Z_{I+1},\dots,Z_{n})$ (hence, $Z\in\mathopen{}\mathclose{{}\left\{Z^{-1},Z^{1}}\right\}$). Algorithm $\mathsf{G}$ computes the following values: [^8] In the proof itself, $\mathsf{G}$ only estimates these values by sampling $R$. For simplicity, here we assume that $\mathsf{G}$ can compute them exactly.
$\mu^{1}={\mathrm{E}}_{R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}\mid R_{I}=1}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z^{1}_{R},W)}\right]$
$\mu^{-1}={\mathrm{E}}_{R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}\mid R_{I}=1}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z^{-1}_{R},W}\right]$
$\mu^{*}={\mathrm{E}}_{R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}\mid R_{I}=0}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z_{R},W)}\right]$
Importantly, we note that when $R_{I}=0$, $Z_{R}$ does not contain $Z_{I}$, and thus $\mathsf{G}$ can compute $\mu^{*}$, given its input. Finally, $\mathsf{G}$ outputs $b\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$ if the following two conditions holds:
$\mathopen{}\mathclose{{}\left|\mu^{b}-\mu^{*}}\right|<1/4k,\ and,\ \mathopen{}\mathclose{{}\left|\mu^{-b}-\mu^{*}}\right|\geq 1/4k.$
Otherwise, $\mathsf{G}$ outputs $\bot$.
Namely, $\mathsf{G}$ outputs $b$ if the expected value of $\mathsf{A}$’s output on $(R,Z_{R})$ when $R_{I}=0$, is closer to the expected value of $\mathsf{A}$’s output on $(R,Z^{b}_{R})$ when $R_{I}=1$ than it is to the expected value of $\mathsf{A}$’s output on $(R,Z^{-b}_{R})$. We next analyze $\mathsf{G}$.
2.3.0.2 Analyzing the predictor.
We will show that the following hold:
  1. 1.
    The value of $R_{I}$ does not change the expectation by much. Formally,
    $\mathopen{}\mathclose{{}\left|{\mathrm{E}}_{R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}\mid R_{I}=1}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z_{R},W)}\right]-{\mathrm{E}}_{R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}\mid R_{I}=0}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z_{R},W)}\right]}\right|<1/4k$
    with probability at least $1-k^{2}/n$, while
  2. 2.
    Filliping the value of $Z_{I}$ does change the expectation. Namely,
    $\mathopen{}\mathclose{{}\left|{\mathrm{E}}_{R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}\mid R_{I}=1}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z^{1}_{R},W)}\right]-{\mathrm{E}}_{R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}\mid R_{I}=1}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z^{-1}_{R},W)}\right]}\right|>1/2k$
    with probability at least $1/k$.
Together, the two items above imply that our distinguisher $\mathsf{G}$ has the properties we need:
  • The first item above implies that $\mathsf{G}$ errs and outputs $-Z_{I}$ with probability at most $k^{2}/n$. Indeed, for $b=-Z_{I}$, it holds that $\mu^{-b}={\mathrm{E}}_{R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}\mid R_{I}=1}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z_{R},W)}\right]$. Thus, Item 1 implies that with all but probability $k^{2}/n$, $\mathopen{}\mathclose{{}\left|\mu^{-b}-\mu^{*}}\right|<1/4k$. In this case, by definition of $\mathsf{G}$, it does not output $b$.
  • In contrast, the two items together imply that $\mathsf{G}$ outputs $Z_{I}$ with probability at least $1/k-k^{2}/n$. Indeed, for $b=Z_{I}$, Item 1 implies that $\mathopen{}\mathclose{{}\left|\mu^{b}-\mu^{*}}\right|<1/4k$ with probability at least $1/k$. Additionally, by the triangular inequality, when both Item 1 and Item 2 hold together we get that
    $\mathopen{}\mathclose{{}\left|{\mathrm{E}}_{R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}\mid R_{I}=0}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z_{R},W)}\right]-{\mathrm{E}}_{R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}\mid R_{I}=1}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z^{-b}_{R},W)}\right]}\right|>1/4k,$
    which by definition implies that $\mathopen{}\mathclose{{}\left|\mu^{-b}-\mu^{*}}\right|\geq 1/4k$ (for $b=Z_{I}$). By a simple union bound, the probability that both Item 1 and Item 2 hold is at least $1/k-k^{2}/n$.
Therefore, assuming that Item 1 and Item 2 hold, this conclude the proof of Section 2.2. We next briefly explain why Item 1 and Item 2 hold with the claimed probability.
2.3.0.3 Bounding the error probability.
We first sketch the proof of Item 1. We claim that
${\mathrm{Pr}}_{Z,W,I}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|{\mathrm{E}}_{R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}\mid R_{I}=0}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z_{R},W)}\right]-{\mathrm{E}}_{R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}\mid R_{I}=1}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z_{R},W)}\right]}\right|\geq 1/4k}\right]\leq k^{2}/n.$ (3)
In the following, we will prove a stronger claim. That is, that for every fixed $Z$ and $W$, it holds that
${\mathrm{Pr}}_{I}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|{\mathrm{E}}_{R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}\mid R_{I}=0}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z_{R},W)}\right]-{\mathrm{E}}_{R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}\mid R_{I}=1}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z_{R},W)}\right]}\right|\geq 1/4k}\right]\leq k^{2}/n.$ (4)
Fix $Z$ and $W$, and let $F(R)=\mathsf{A}(R,Z_{R},W)$. Using this notation, we can rewrite Equation 4 as
${\mathrm{Pr}}_{I}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|{\mathrm{E}}_{R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}\mid R_{I}=0}\mathopen{}\mathclose{{}\left[F(R)}\right]-{\mathrm{E}}_{R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}\mid R_{I}=1}\mathopen{}\mathclose{{}\left[F(R)}\right]}\right|\geq 1/4k}\right]\leq k^{2}/n.$ (5)
Finally, Equation 5 follows a more general well-known fact on the uniform distribution over random strings: For a random string $R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}$, and a random index $I\leftarrow[n]$, the statistical distance between $R|_{R_{I}=0}$ and $R|_{R_{I}=1}$ is at most $1/\sqrt{n}$. This fact implies that the function $F$ cannot distinguish too well between samples from $R|_{R_{I}=0}$ and samples from $R|_{R_{I}=1}$. [^9] Importantly, notice that $F$ does not get $I$ as input. We prove the following stronger lemma that finishes this part of the proof.
[…]
Let $R\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}$. For any (randomized) function $F:\{0,1\}^{n}\rightarrow\{0,1\}$ and any $k>0$, it holds that
${\mathrm{Pr}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|\>{\mathrm{E}}\mathopen{}\mathclose{{}\left[F(R)\mid R_{i}=0}\right]-{\mathrm{E}}\mathopen{}\mathclose{{}\left[F(R)\mid R_{i}=1}\right]\>}\right|\geq 1/4k}\right]\leq\Theta\mathopen{}\mathclose{{}\left(\frac{k^{2}}{n}}\right),$ (6)
where the expectations are taken over $R$ and the randomness of $F$.
2.3.0.4 Lower bounding the distinguishing advantage.
Next we sketch the proof of Item 2. We need to show that
${\mathrm{Pr}}_{Z,W,I}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|{\mathrm{E}}_{R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}\mid R_{I}=1}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z^{1}_{R},W)}\right]-{\mathrm{E}}_{R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}\mid R_{I}=1}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z^{-1}_{R},W)}\right]}\right|\geq 1/2k}\right]\geq 1/k.$ (7)
Or equivalently,
${\mathrm{Pr}}_{Z,W,I}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|{\mathrm{E}}_{R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}\mid R_{I}=1}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z_{R},W)}\right]-{\mathrm{E}}_{R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}\mid R_{I}=1}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z^{(I)}_{R},W)}\right]}\right|\geq 1/2k}\right]\geq 1/k.$ (8)
We claim that Equation 8 follows by the assumption that $\mathsf{A}$ is a good distinguisher. Indeed, assume, for the sake of contradiction, that this is not the case. Then, by a simple computation (and using the fact that $\mathsf{A}(R,Z_{R},W)\in\mathopen{}\mathclose{{}\left\{0,1}\right\}$) we get that
${\mathrm{E}}_{Z,W,I}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|{\mathrm{E}}_{R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}\mid R_{I}=1}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z_{R},W)}\right]-{\mathrm{E}}_{R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}\mid R_{I}=1}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z^{(I)}_{R},W)}\right]}\right|}\right]\leq 2/k.$
and thus, using the triangular inequality, we can write,
$\mathopen{}\mathclose{{}\left|{\mathrm{E}}_{Z,W,I}\mathopen{}\mathclose{{}\left[{\mathrm{E}}_{R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}\mid R_{I}=1}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z_{R},W)}\right]-{\mathrm{E}}_{R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}\mid R_{I}=1}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z^{(I)}_{R},W)}\right]}\right]}\right|<2/k.$
Equivalently, we have that,
$\mathopen{}\mathclose{{}\left|{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z_{R},W)=1\mid R_{I}=1}\right]-{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z^{(I)}_{R},W)=1\mid R_{I}=1}\right]}\right|<2/k.$
which means that $\mathsf{A}$ fails to distinguish between $(R,Z_{R})$ and $(R,Z^{(I)}_{R})$ when $R_{I}=1$. On the other hand, we notice that $\mathsf{A}$ also must fail when $R_{I}=0$. Indeed, when $R_{I}=0$ it holds that $Z_{R}=Z^{(I)}_{R}$, and thus we get that
$\mathopen{}\mathclose{{}\left|{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z_{R},W)=1\mid R_{I}=0}\right]-{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z^{(I)}_{R},W)=1\mid R_{I}=0}\right]}\right|=0.$
Finally, since $\mathsf{A}$ fails to distinguish $Z_{R}$ from $Z^{(I)}_{R}$ both when $R_{I}=0$ and $R_{I}=1$, we conclude that $\mathsf{A}$ is not a good distinguisher. That is,
$\mathopen{}\mathclose{{}\left|{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z_{R},W)=1}\right]-{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{A}(R,Z^{(I)}_{R},W)=1}\right]}\right|<1/k,$
which contradicts our assumption on $\mathsf{A}$.
3 Related Works
Computational differential privacy. Computational differential privacy (${\sf CDP}$) can be defined using two primary approaches. The more flexible and widely used approach is the indistinguishability-based definition, which limits the distinguished $g$ in Section 1, to those that are computationally efficient. The second approach, known as the simulation-based definition, requires that the output distribution of the mechanism $f$ be computationally indistinguishable from that of an (information-theoretic) differentially private mechanism. Various relationships between these and other privacy definitions have been established in [MPRV09] (a more updated survey is provided in [MeisingsethR25]). We remark that our result holds for the weaker definition of indistinguishability-based ${\sf CDP}$, which makes it stronger. As a corollary, we also get the equivalence of the power of the definitions for the inner-product task we consider (in our accuracy regime).
CDP in the centralized model. In the single-party scenario (i.e., the centralized model), computational and information-theoretic differential privacy appear to be more closely aligned. Specifically, [GKY11] demonstrated that a broad class of ${\sf CDP}$ mechanisms can be converted into an information-theoretic ${\sf DP}$ mechanism. On the other hand, [BunCV16, GhaziIK0M23] showed that under certain non-standard and very-strong cryptographic assumptions, there exist somewhat contrived tasks that can be efficiently solved with ${\sf CDP}$, but remain infeasible ([BunCV16]) or impossible ([GhaziIK0M23]) under information-theoretic ${\sf DP}$. It still remains open whether such separations exist under more standard cryptographic assumptions and for more natural tasks.
CDP in the local model. At the other end of the spectrum, the local model is highly relevant in practical applications. In this setting, each of the (typically many) participants holds a single data element. Protocols achieving information-theoretic ${\sf DP}$ in this model often rely on randomized response, which has been shown to be optimal for counting functions, including inner product, as proven by [chan2012optimal]. In contrast, local ${\sf CDP}$ protocols can leverage secure multiparty computation to simulate any efficient single-party mechanism, thereby demonstrating a fundamental gap between the power of ${\sf CDP}$ and information-theoretic ${\sf DP}$. We remark that there exist other approaches to bridge this by relaxing the distributed model, e.g., using a trusted shuffler [BittauEMMRLRKTS17, CheuSUZZ19] (see [CheuSurvey22] for a survey on this model), or partially trusted servers that enable more accurate estimations using weaker and more practical cryptographic tools than secure multi-party computation (e.g., [BangaloreCV24, DamgardKNOP24, BohlerK20]).
Two-Party CDP. In the two-party (or small-party) setting, the complexity of ${\sf CDP}$ protocols is much less clear. Prior works have maintly focused on Boolean functionalities, where each party holds a single sensitive bit, and the objective is to compute a Boolean function over these bits while preserving privacy (e.g., XOR). [GMPS13] established that, for any non-trivial Boolean functionality, there is a fundamental gap in accuracy between what can be achieved in the centralized and distributed settings. Moreover, any ${\sf CDP}$ protocol that surpasses this accuracy gap would imply the existence of one-way functions. Later, [GKMPS16] demonstrated that an accurate enough ${\sf CDP}$ protocol for the XOR function would inherently imply an oblivious transfer protocol. Building on this, [HNOSS20] proved that any meaningful $\varepsilon$-CDP two-party protocol for XOR necessarily implies an (infinitely-often) key agreement protocol. [HMSS19] refined and extended the results of [GKMPS16, HNOSS20], showing that any non-trivial ${\sf CDP}$ protocol for XOR implies oblivious transfer.
Beyond Boolean functionalities, the complexity of ${\sf CDP}$ protocols for more general tasks such as computing low-sensitivity many-bit functions like the inner product remains largely unexplored. The only exceptions are the work of [haitner2016limits], who applied a generic reduction to the impossibility result of [MPRV09], concluding that no accurate ${\sf CDP}$ protocol for inner product can exist in the random oracle model, and more recently, the work of [HaitnerMST22] who showed that any non-trivial ${\sf CDP}$ inner-product protocol implies a key-agreement protocol. But none of these results close the gap to ${\sf OT}$, and thus, our result provide the first tight characterization (with respect to assumptions) of a natural, non-boolean functionality.
4 Preliminaries
4.1 Distributions and Random Variables
The support of a distribution $P$ over a finite set ${\cal S}$ is defined by $\operatorname{Supp}(P):=\mathopen{}\mathclose{{}\left\{x\in{\cal S}:P(x)>0}\right\}$. For a distribution or a random variable $D$, let $d\leftarrow D$ denote that $d$ was sampled according to $D$. Similarly, for a set ${\cal S}$, let $x\leftarrow{\cal S}$ denote that $x$ is drawn uniformly from ${\cal S}$, and denote by ${\cal U}_{{\cal S}}$ the uniform distribution over ${\cal S}$. For a finite set ${\cal X}$ and a distribution $C_{X}$ over ${\cal X}$, we use the capital letter $X$ to denote the random variable that takes values in ${\cal X}$ and is sampled according to $C_{X}$. The statistical distance (also known as, variation distance) of two distributions $P$ and $Q$ over a discrete domain ${\cal X}$ is defined by $\mathsf{\textsc{SD}}\mathopen{}\mathclose{{}\left(P,Q}\right):=\max_{{\cal S}\subseteq{\cal X}}\mathopen{}\mathclose{{}\left|P({\cal S})-Q({\cal S})}\right|=\frac{1}{2}\sum_{x\in{\cal S}}\mathopen{}\mathclose{{}\left|P(x)-Q(x)}\right|$. For a vector $x=(x_{1},\ldots,x_{n})$ and index $i\in[n]$, we let $x_{-i}=(x_{1},\ldots,x_{i-1},x_{i+1},\ldots,x_{n})$ and $x^{(i)}=(x_{1},\ldots,x_{i-1},-x_{i},x_{i+1},\ldots,x_{n})$, for a set ${\cal S}\subseteq[n]$ we let $x_{{\cal S}}=(x_{i})_{i\in{\cal S}}$ and $x_{-{\cal S}}=(x_{i})_{i\in[n]\setminus{\cal S}}$, and for a vector $r\in\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}$ we let $x_{r}=x_{\mathopen{}\mathclose{{}\left\{i\colon r_{i}=1}\right\}}$ and $x_{-r}=x_{\mathopen{}\mathclose{{}\left\{i\colon r_{i}=0}\right\}}$.
[…]
4.2 Differential Privacy
[…]
A randomized function $f\colon\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}\mapsto{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{\ast}}$ is an $n$-size, $(\varepsilon,\delta)$-differentially private mechanism (denoted $(\varepsilon,\delta)$-DP) if for every neighboring $w,w^{\prime}\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ and every function $g\colon{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{\ast}}\mapsto\mathopen{}\mathclose{{}\left\{0,1}\right\}$, it holds that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[g(f(w))=1}\right]\leq e^{\varepsilon}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[g(f(w^{\prime}))=1}\right]+\delta.$
If $\delta=0$, we omit it from the notation.
4.2.1 Computational Differential Privacy
There are several ways for defining computational differential privacy (see Section 3). We use the most relaxed version due to [BNO08]. For notational convenience, we focus on databases over $\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$.
[…]
A randomized function ensemble $f=\mathopen{}\mathclose{{}\left\{f_{\kappa}\colon\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n(\kappa)}\mapsto{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{\ast}}}\right\}$ is an $n$-size, $(\varepsilon,\delta)$-computationally differentially private (denoted $(\varepsilon,\delta)$-${\sf CDP}$) if for every poly-size circuit family $\mathopen{}\mathclose{{}\left\{\mathsf{A}_{\kappa}}\right\}_{\kappa\in{\mathbb{N}}}$, the following holds for every large enough $\kappa$ and every neighboring $w,w^{\prime}\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n(\kappa)}$:
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{A}_{\kappa}(f_{\kappa}(w))=1}\right]\leq e^{\varepsilon(\kappa)}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{A}_{\kappa}(f_{\kappa}(w^{\prime}))=1}\right]+\delta(\kappa).$
If $\delta(\kappa)=\operatorname{neg}(\kappa)$, we omit it from the notation.
4.2.2 Two-Party Differential Privacy
In this section we formally define distributed differential privacy mechanism (i.e., protocols).
[…]
A two-party protocol $\Pi=(\mathsf{A},\mathsf{B})$ is $(\varepsilon,\delta)$-differentially private, denoted $(\varepsilon,\delta)$-${\sf DP}$, if the following holds for every algorithm $\mathsf{D}$: let $\operatorname{V}^{\mathsf{P}}(x,y)(\kappa)$ be the view of party $\mathsf{P}$ in a random execution of $\Pi(x,y)(1^{\kappa})$. Then for every $\kappa,n\in{\mathbb{N}}$, $x\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ and neighboring $y,y^{\prime}\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$:
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{D}(V^{\mathsf{A}}(x,y)(\kappa))=1}\right]\leq e^{\varepsilon(\kappa)}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{D}(V^{\mathsf{A}}(x,y^{\prime})(\kappa))=1}\right]+\delta(\kappa),$
and for every $y\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ and neighboring $x,x^{\prime}\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$:
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{D}(V^{\mathsf{B}}(x,y)(\kappa))=1}\right]\leq e^{\varepsilon(\kappa)}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{D}(V^{\mathsf{B}}(x^{\prime},y)(\kappa))=1}\right]+\delta(\kappa).$
Protocol $\Pi$ is $(\varepsilon,\delta)$-computational differentially private, denoted $(\varepsilon,\delta)$-${\sf CDP}$, if the above inequalities only hold for a non-uniform ppt $\mathsf{D}$ and large enough $\kappa$. We omit $\delta=\operatorname{neg}(\kappa)$ from the notation.  [^10] Note that define we give for two-party differentially private protocols is a semi-honest definition, in which we ask for the security to hold when the parties interact in an honest execution of the protocol. Since we are proving a lower bound, starting from this weaker guarantee (as opposed to security against malicious players), yields a stronger result.
[…]
An alternative, stronger definition of computational differential privacy, known as simulation-based computational differential privacy, requires that the distribution of each party’s view be computationally indistinguishable from a distribution that ensures privacy in an information-theoretic sense. Section 4.2.2 is a weaker notion in comparison. Consequently, establishing a lower bound for a protocol that satisfies this weaker guarantee (as we do in this work) yields a stronger result.
4.3 Useful Claims
The following two propositions state that given the output of a differentially private function, it is not possible to predict well even a random index (even if all other indexes are leaked). The first proposition handles the information-theoretic case and the second handles the computation case. Both propositions are proven in Section C.1.
[…]
Let $f\colon\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}\rightarrow{\cal Y}$ be an $(\varepsilon,\delta)$-DP function, let $g\colon[n]\times\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n-1}\times{\cal Y}\rightarrow\mathopen{}\mathclose{{}\left\{-1,1,\bot}\right\}$ be a (randomized) function, and let $X=(X_{1},\ldots,X_{n})\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$. Then the following holds for every $i\in[n]$ where $X_{i}^{*}=g(i,X_{-i},f(X_{1},\ldots,X_{n}))$:
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X_{i}^{*}=X_{i}}\right]\leq e^{\varepsilon}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X_{i}^{*}=-X_{i}}\right]+\delta.$
[…]
Let $f=\mathopen{}\mathclose{{}\left\{f_{\kappa}\colon\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n(\kappa)}\rightarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{m(\kappa)}}\right\}_{\kappa\in{\mathbb{N}}}$ be an $(\varepsilon,\delta)$-CDP function ensemble, and let $\mathopen{}\mathclose{{}\left\{g_{\kappa}}\right\}_{\kappa\in{\mathbb{N}}}$ be a poly-size circuit family. Then, for large enough $\kappa$ and $X=(X_{1},\ldots,X_{n(\kappa)})\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n(\kappa)}$, the following holds for every $i\in[n(\kappa)]$ where $X_{i}^{*}=g_{\kappa}(i,X_{-i},f_{\kappa}(X_{1},\ldots,X_{n}))$:
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X_{i}^{*}=X_{i}}\right]\leq e^{\varepsilon(\kappa)}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X_{i}^{*}=-X_{i}}\right]+\delta(\kappa).$
[…]
Let $R\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}$. For any function $f:\{0,1\}^{n}\rightarrow\{0,1\}$ and $\alpha<0.01$, it holds that
${\mathrm{Pr}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[\>\mathopen{}\mathclose{{}\left|\>\mathbb{E}[f(R)\mid R_{i}=0]-\mathbb{E}[f(R)\mid R_{i}=1]\>}\right|\geq\alpha}\right]\leq\frac{2+2\log(\frac{1}{\alpha})}{n\alpha^{2}}.$
Proof.
Define $S_{1}=\{r\in\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\colon f(r)=1\}$. Then for any $i\in[n]$, we have
$\begin{array}[]{rl}\mathopen{}\mathclose{{}\left|\mathbb{E}[f(R)\mid R_{i}=0]-\mathbb{E}[f(R)\mid R_{i}=1]}\right|&=\mathopen{}\mathclose{{}\left|{\mathrm{Pr}}[R\in S_{1}|R_{i}=0]-{\mathrm{Pr}}[R\in S_{1}|R_{i}=1]}\right|\\ &=\mathopen{}\mathclose{{}\left|\frac{{\mathrm{Pr}}[R_{i}=0|R\in S_{1}]\cdot{\mathrm{Pr}}[R\in S_{1}]}{{\mathrm{Pr}}[R_{i}=0]}-\frac{{\mathrm{Pr}}[R_{i}=1|R\in S_{1}]\cdot{\mathrm{Pr}}[R\in S_{1}]}{{\mathrm{Pr}}[R_{i}=1]}}\right|\\ &=\frac{2\mathopen{}\mathclose{{}\left|S_{1}}\right|}{2^{n}}\mathopen{}\mathclose{{}\left|{\mathrm{Pr}}[R_{i}=0|R\in S_{1}]-{\mathrm{Pr}}[R_{i}=1|R\in S_{1}]}\right|\end{array}$
When $|S_{1}|\leq\alpha\cdot 2^{n-1}$, we have $\mathopen{}\mathclose{{}\left|\mathbb{E}[f(R)\mid R_{i}=0]-\mathbb{E}[f(R)\mid R_{i}=1]}\right|\leq\frac{2\mathopen{}\mathclose{{}\left|S_{1}}\right|}{2^{n}}\leq\alpha$ for any $i\in[n]$. Hence, in the following, we assume $|S_{1}|>\alpha\cdot 2^{n-1}$.
Define $I_{bad}=\{i\colon\mathopen{}\mathclose{{}\left|{\mathrm{Pr}}[R_{i}=0|R\in S_{1}]-{\mathrm{Pr}}[R_{i}=1|R\in S_{1}]}\right|\geq 2\alpha\}$ and $k=\mathopen{}\mathclose{{}\left|I_{bad}}\right|$, and denote $I_{bad}=\{i_{1},\dots,i_{k}\}$. Define $(X_{i_{1}},\ldots X_{i_{k}})=(R_{i_{1}},\dots,R_{i_{k}})\mid_{R\in S_{1}}$. Consider the min-entropy
$\begin{array}[]{rl}H_{min}(X_{i_{1}},\dots,X_{i_{k}})&\leq H(X_{i_{1}},\dots,X_{i_{k}})\\ &\leq\sum_{j=1}^{k}H(X_{i_{j}})\\ &\leq k\cdot\mathopen{}\mathclose{{}\left(-(\frac{1}{2}+2\alpha)\cdot\log(\frac{1}{2}+2\alpha)-(\frac{1}{2}-2\alpha)\cdot\log(\frac{1}{2}-2\alpha)}\right)\\ &=k\cdot\mathopen{}\mathclose{{}\left(-(\frac{1}{2}+2\alpha)\cdot(\log(1+4\alpha)-1)-(\frac{1}{2}-2\alpha)\cdot(\log(1-4\alpha)-1)}\right)\\ &=k\cdot\mathopen{}\mathclose{{}\left(1-(\frac{1}{2}+2\alpha)\cdot\log(1+4\alpha)-(\frac{1}{2}-2\alpha)\cdot\log(1-4\alpha)}\right),\par\end{array}$
where $H_{min}(Y)$ is the minimum entropy of $Y$ and $H(Y)$ is the Shannon entropy of $Y$. The third inequality holds since by the definition of $I_{bad}$, for every $j\in[k]$ it holds that $\mathopen{}\mathclose{{}\left|{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X_{i_{j}}=1}\right]-{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X_{i_{j}}=0}\right]}\right|>2\alpha$, and therefore $H(X_{i_{j}})\leq H(1/2+2\alpha)$.
Therefore, there exists $b_{1},\dots,b_{k}\in\{0,1\}$, such that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[(R_{i_{1}},\ldots,R_{i_{k}})=(b_{1},\ldots,b_{k})\mid R\in S_{1}}\right]$ $={\mathrm{Pr}}\mathopen{}\mathclose{{}\left[(X_{i_{1}},\ldots,X_{i_{k}})=(b_{1},\ldots,b_{k})}\right]$ (9)
$=2^{-H_{min}(X_{i_{1}},\dots,X_{i_{k}})}$
$\geq 2^{k\cdot\mathopen{}\mathclose{{}\left(-1+(\frac{1}{2}+2\alpha)\cdot\log(1+4\alpha)+(\frac{1}{2}-2\alpha)\cdot\log(1-4\alpha)}\right)}.$
Let $S_{bad}=\{r\in\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\colon\mathopen{}\mathclose{{}\left\{(r_{i_{1}},\ldots,r_{i_{k}})=(b_{1},\ldots,b_{k})}\right\}\land\mathopen{}\mathclose{{}\left\{r\in S_{1}}\right\}\}$. It holds that
$|S_{bad}|$ $=\mathopen{}\mathclose{{}\left|S_{1}}\right|\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[(R_{i_{1}},\ldots,R_{i_{k}})=(b_{1},\ldots,b_{k})\mid R\in S_{1}}\right]$
$\geq\alpha\cdot 2^{n-1}\cdot 2^{k\cdot\mathopen{}\mathclose{{}\left(-1+(\frac{1}{2}+2\alpha)\cdot\log(1+4\alpha)+(\frac{1}{2}-2\alpha)\cdot\log(1-4\alpha)}\right)},$
where the inequality holds by Equation 9 and since $\mathopen{}\mathclose{{}\left|S_{1}}\right|\geq\alpha\cdot 2^{n-1}$. Notice that any string in $S_{bad}$ depends on at most $n-k$ bits. It implies that $|S_{bad}|\leq 2^{n-k}$. Therefore, we have
$\begin{array}[]{rl}&2^{n-k}\geq\alpha\cdot 2^{n-1}\cdot 2^{k\cdot\mathopen{}\mathclose{{}\left(-1+(\frac{1}{2}+2\alpha)\cdot\log(1+4\alpha)+(\frac{1}{2}-2\alpha)\cdot\log(1-4\alpha)}\right)}\\ \Rightarrow&n-k\geq\log\alpha+n-1+k\cdot\mathopen{}\mathclose{{}\left(-1+(\frac{1}{2}+2\alpha)\cdot\log(1+4\alpha)+(\frac{1}{2}-2\alpha)\cdot\log(1-4\alpha)}\right)\\ \Rightarrow&1-\log\alpha\geq k\cdot((\frac{1}{2}+2\alpha)\cdot\log(1+4\alpha)+(\frac{1}{2}-2\alpha)\cdot\log(1-4\alpha))\\ \Rightarrow&1-\log\alpha\geq k\cdot(4\alpha\cdot\log(1+4\alpha)+(\frac{1}{2}-2\alpha)\cdot\log(1-16\alpha^{2}))\\ \Rightarrow&1-\log\alpha\geq k\cdot(15.9\alpha^{2}-8\alpha^{2}+32\alpha^{3})=k\cdot(7.9\alpha^{2}+32\alpha^{3})>0.5k\alpha^{2}\\ \Rightarrow&k\leq\frac{2-2\log\alpha}{\alpha^{2}}=\frac{2+2\log(1/\alpha)}{\alpha^{2}},\end{array}$
Where the third transition holds since
$(\frac{1}{2}+2\alpha)\cdot\log(1+4\alpha)+(\frac{1}{2}-2\alpha)\cdot\log(1-4\alpha)$
$=4\alpha\cdot\log(1+4\alpha)+(\frac{1}{2}-2\alpha)\mathopen{}\mathclose{{}\left(\log(1+4\alpha)+\log(1-4\alpha)}\right)$
$=4\alpha\cdot\log(1+4\alpha)+(\frac{1}{2}-2\alpha)\cdot\log(1-16\alpha^{2}),$
and the forth transition holds since $4\alpha\cdot\log(1+4\alpha)+(\frac{1}{2}-2\alpha)\cdot\log(1-16\alpha^{2})>15.9\alpha^{2}-8\alpha^{2}+32\alpha^{3}$ for $\alpha<0.01$. Thus, we conclude that
${\mathrm{Pr}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|\mathbb{E}[f(R)\mid R_{i}=0]-\mathbb{E}[f(R)\mid R_{i}=1]}\right|\geq\alpha}\right]\leq\frac{k}{n}\leq\frac{2+2\log(1/\alpha)}{n\alpha^{2}}.$
$\Box$
4.4 Channels and Two-Party Protocols
4.4.0.1 Channels.
A channel is simply a distribution of a pair of tuples defined as follows.
[…]
A channel $C_{(X,U)(Y,V)}$ of size $n$ over alphabet $\Sigma$ is a probability distribution over $(\Sigma^{n}\times\mathopen{}\mathclose{{}\left\{0,1}\right\}^{\ast})\times(\Sigma^{n}\times\mathopen{}\mathclose{{}\left\{0,1}\right\}^{\ast})$. The ensemble $C_{(X,U)(Y,V)}=\mathopen{}\mathclose{{}\left\{C_{(X_{\kappa},U_{\kappa})(Y_{\kappa},V_{\kappa})}}\right\}_{\kappa\in{\mathbb{N}}}$ is an $n$-size channel ensemble, if for every $\kappa\in{\mathbb{N}}$, $C_{(X_{\kappa},U_{\kappa})(Y_{\kappa},V_{\kappa})}$ is an $n(\kappa)$-size channel. We refer to $X$ and $Y$ as the local outputs, and to $U$ and $V$ as the views.
We view a channel as the experiment in which there are two parties $\mathsf{A}$ and $\mathsf{B}$. Party $\mathsf{A}$ receives “output” $X$ and “view” $U$, and party $\mathsf{B}$ receives “output” $Y$ and “view” $V$. Unless stated otherwise, the channels we consider are over the alphabet $\Sigma=\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$. We naturally identify channels with the distribution that characterizes their output.
4.4.1 Two-Party Protocols
A two-party protocol $\Pi=(\mathsf{A},\mathsf{B})$ is ppt if the running time of both parties is polynomial in their input length. We let $\Pi(x,y)(z)$ or $(\mathsf{A}(x),\mathsf{B}(y))(z)$ denote a random execution of $\Pi$ on a common input $z$, and private inputs $x,y$.
[…]
In a two-party protocol $\Pi$ with oracle access to a protocol $\Psi$, denoted $\Pi^{\Psi}$, the parties make use of the next-message function of $\Psi$ [^11] The function that on a partial view of one of the parties, returns its next message. In a two-party protocol $\Pi$ with oracle access to a channel $C_{ZW}$, denoted $\Pi^{C}$, the parties can jointly invoke $C$ for several times. In each call, an independent pair $(z,w)$ is sampled according to $C_{ZW}$, one party gets $z$, the other gets $w$.
[…]
For a two-party protocol $\Pi$ that gets a security parameter $1^{\kappa}$ as its (only, common) input, we associate the channel ensemble $\mathopen{}\mathclose{{}\left\{C_{\Pi(1^{\kappa})}}\right\}_{\kappa\in{\mathbb{N}}}$.
[…]
A channel $C=C_{(X,U),(Y,V)}$ is $(\alpha,\gamma)$-accurate for the function $f$, if ${\mathrm{Pr}}_{C}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|{\rm out}(V)-f(X,Y)}\right|\leq\alpha}\right]\geq\gamma$, where ${\rm out}(V)$ is the designated output. A channel ensemble $C_{(X,U),(Y,V)}=\mathopen{}\mathclose{{}\left\{C_{(X_{\kappa},U_{\kappa}),(Y_{\kappa},V_{\kappa})}}\right\}_{\kappa\in{\mathbb{N}}}$ is $(\alpha,\gamma)$-accurate for $f$ if $C_{(X_{\kappa},U_{\kappa}),(Y_{\kappa},V_{\kappa})}$ is $(\alpha(\kappa),\gamma(\kappa))$-accurate for $f$, for every $\kappa\in{\mathbb{N}}$.
4.4.2 Differentially Private Channels
Differentially private channels are naturally defined as follows:
[…]
An $n$-size channel $C=C_{(X,U),(Y,V)}$ with $X,Y$ over $\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ is $(\varepsilon,\delta)$-differentially private (denoted $(\varepsilon,\delta)$-${\sf DP}$) if for every $x\in\operatorname{Supp}(X)$ there exists an $n$-size $(\varepsilon,\delta)$-${\sf DP}$ mechanisms $\mathsf{M}_{x}$ such that $(X,Y,U)\equiv(X,Y,\mathsf{M}_{X}(Y))$, and for every $y\in\operatorname{Supp}(Y)$ there exists an $n$-size $(\varepsilon,\delta)$-${\sf DP}$ mechanisms $\mathsf{M}_{y}^{\prime}$ such that $(X,Y,V)\equiv(X,Y,\mathsf{M}_{Y}^{\prime}(X))$. In addition, we say that the channel is uniform if $X$ and $Y$ are independent random variables uniformly distributed in $\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$.
[…]
An $n$-size channel ensemble $C=\mathopen{}\mathclose{{}\left\{C_{(X_{\kappa},U_{\kappa}),(Y_{\kappa},V_{\kappa})}}\right\}_{\kappa\in{\mathbb{N}}}$ with $X_{\kappa},Y_{\kappa}$ over $\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ is $(\varepsilon,\delta)$-computationally differentially private (denoted $(\varepsilon,\delta)$-${\sf CDP}$) if for every ensemble $\mathopen{}\mathclose{{}\left\{x_{\kappa}\in\operatorname{Supp}(X_{\kappa})}\right\}_{\kappa\in{\mathbb{N}}}$ there exists an $n$-size $(\varepsilon,\delta)$-CDP mechanisms ensemble $\mathopen{}\mathclose{{}\left\{\mathsf{M}_{x_{\kappa}}}\right\}_{\kappa\in{\mathbb{N}}}$ such that $(X_{\kappa},Y_{\kappa},U_{\kappa})\equiv(X_{\kappa},Y_{\kappa},\mathsf{M}_{X_{\kappa}}(Y_{\kappa}))$, for every $\kappa\in{\mathbb{N}}$, and for every ensemble $\mathopen{}\mathclose{{}\left\{y_{\kappa}\in\operatorname{Supp}(Y_{\kappa})}\right\}_{\kappa\in{\mathbb{N}}}$ there exists an $n$-size $(\varepsilon,\delta)$-${\sf CDP}$ mechanisms ensemble $\mathopen{}\mathclose{{}\left\{\mathsf{M}^{\prime}_{y_{\kappa}}}\right\}_{\kappa\in{\mathbb{N}}}$ such that $(X_{\kappa},Y_{\kappa},V_{\kappa})\equiv(X_{\kappa},Y_{\kappa},\mathsf{M}_{Y_{\kappa}}^{\prime}(X_{\kappa}))$ for every $\kappa\in{\mathbb{N}}$. In addition, we say that the channel is uniform if $X_{\kappa}$ and $Y_{\kappa}$ are independent random variables uniformly distributed in {±$\{\pm 1\}^{n}$ for all $\kappa\in{\mathbb{N}}$.
4.4.3 Weak Erasure Channel (WEC)
[…]
A channel $((O_{A},V_{A}),(O_{B},V_{B}))$ with $O_{A}\in\mathopen{}\mathclose{{}\left\{0,1}\right\}$ and $O_{B}\in\mathopen{}\mathclose{{}\left\{0,1,\bot}\right\}$ is a weak erasure channel, denoted $(\alpha,p,q)$-${\sf WEC}$, if:
  • Random erasure: ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[O_{B}=\perp}\right]=1/2$.
  • Agreement: ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[O_{A}\neq O_{B}\mid O_{B}\neq\bot}\right]\leq\alpha$.
  • Secrecy:
    1. 1.
      For every algorithm $\mathsf{D}$ it holds that
      $\mathopen{}\mathclose{{}\left|{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{D}(V_{A})=1\mid O_{B}\neq\perp}\right]-{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{D}(V_{A})=1\mid O_{B}=\perp}\right]}\right|\leq p$
      (Alice doesn’t know if $O_{B}=\perp$.)
    2. 2.
      For every algorithm $\mathsf{D}$ it holds that
      ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{D}(V_{B})=O_{A}\mid O_{B}=\bot}\right]\leq\frac{1+q}{2}.$
      (i.e., if $O_{B}=\bot$, Bob don’t know what is the value of $O_{A}$).
We say that a channel ensemble $C=\mathopen{}\mathclose{{}\left\{C_{\kappa}}\right\}_{\kappa\in N}$ is a computational weak erasure channel, denoted $(\alpha,p,q)$-CompWEC, if for every ppt algorithm $\mathsf{D}$ and every sufficiently large $\kappa\in{\mathbb{N}}$, $C_{\kappa}$ satisfies the properties stated in the items above, where the secrecy property holds with respect to a ppt algorithm $\mathsf{D}$. A protocol $\Lambda$ is said to be $(\alpha,p,q)$-${\sf CompWEC}$, if the ensemble induces by the protocol (that is, $C=\mathopen{}\mathclose{{}\left\{C_{\Lambda(\kappa)}}\right\}_{\kappa\in{\mathbb{N}}}$) is $(\alpha,p,q)$-${\sf CompWEC}$.
4.4.4 Approximate Weak Erasure Channel (AWEC)
[…]
A channel $C=((O_{A},V_{A}),(O_{B},V_{B}))$ over $([-n,n]\times\mathopen{}\mathclose{{}\left\{0,1}\right\}^{*})\times(([-n,n]\cup\bot)\times\mathopen{}\mathclose{{}\left\{0,1}\right\}^{*})$ is an approximate weak erasure channel, denoted $(\ell,\alpha,p,q)$-AWEC if:
  • Random erasure: ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[O_{B}=\perp}\right]=1/2$.
  • Accuracy: ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|O_{A}-O_{B}}\right|>\ell\mid O_{B}\neq\bot}\right]\leq\alpha$.
  • Secrecy:
    1. 1.
      For every algorithm $\mathsf{D}$ it holds that
      $\mathopen{}\mathclose{{}\left|{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{D}(V_{A})=1\mid O_{B}\neq\perp}\right]-{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{D}(V_{A})=1\mid O_{B}=\perp}\right]}\right|\leq p$
      (Alice doesn’t know if $O_{B}=\bot$).
    2. 2.
      For every algorithm $\mathsf{D}$ it holds that
      ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|\mathsf{D}(V_{B})-O_{A}}\right|\leq 1000\ell\mid O_{B}=\bot}\right]\leq q.$
      (i.e., if $O_{B}=\bot$, Bob can’t estimate the value of $O_{A}$ with error $\leq 1000\ell$).
We say that a channel ensemble $C=\mathopen{}\mathclose{{}\left\{C_{\kappa}}\right\}_{\kappa\in N}$ is a computational approximate weak erasure channel, denoted $(\ell,\alpha,p,q)$-CompAWEC, if for every ppt algorithm $\mathsf{D}$ and every sufficiently large $\kappa\in{\mathbb{N}}$, $C_{\kappa}$ satisfies the properties stated in the items above. A protocol $\Gamma$ is said to be $(\ell,\alpha,p,q)$-${\sf CompAWEC}$, if the ensemble induced by the protocol (that is, $C=\mathopen{}\mathclose{{}\left\{C_{\Gamma(\kappa)}}\right\}_{\kappa\in{\mathbb{N}}}$) is $(\ell,\alpha,p,q)$-${\sf CompAWEC}$.
We will make use of the following lemma, which shows that for some choices of the parameters, AWEC implies WEC. The lemma is proven in Appendix B.
[…]
For every $\ell>0$, there exists a ppt protocol $\Lambda=(\mathsf{P}_{1},\mathsf{P}_{2})$ such that given an oracle access to an $(\ell,\alpha,p,q)$-AWEC $C$, the channel $\tilde{C}$ induced by $\Lambda^{C}$ is $(\alpha^{\prime}=\alpha+0.001,\>p^{\prime}=p,\>q^{\prime}=1/2+2(q+0.01))$-WEC. Furthermore, the proof is constructive in a black-box manner:
  1. 1.
    There exists an oracle-aided ppt algorithm $\mathsf{E}_{1}$ such that for every channel $C=((O_{\mathsf{A}},{V_{\mathsf{A}}}),(O_{\mathsf{B}},{V_{\mathsf{B}}}))$ and algorithm $\mathsf{D}$ violating the WEC secrecy property 1 of $\tilde{C}$, algorithm $\mathsf{E}_{1}^{\mathsf{D}}$ violates the AWEC secrecy property 1 of $C$.
  2. 2.
    There exists an oracle-aided ppt algorithm $\mathsf{E}_{2}$ such that for every channel $C=((O_{\mathsf{A}},{V_{\mathsf{A}}}),(O_{\mathsf{B}},{V_{\mathsf{B}}}))$ and algorithm $\mathsf{D}$ violating the WEC secrecy property 2 of $\tilde{C}$, algorithm $\mathsf{E}_{2}^{\mathsf{D}}$ violates the AWEC secrecy property 2 of $C$.
Since Section 4.4.4 is constructive, the following is an immediate corollary.
[…]
There exists an oracle aided ppt protocol $\Lambda$, such that given a protocol $\Gamma$ that induces $(\ell,\alpha,p,q)$-CompAWEC, it holds that $\Lambda^{\Gamma}$ is $(\alpha^{\prime}=\alpha+0.001,\>p^{\prime}=p,\>q^{\prime}=1/2+2(q+0.01))$-CompWEC.
Proof of 4.4.4.
Let $\Lambda$ be the ppt algorithm guaranteed by Lemma 4.4.4. Given an $(\ell,\alpha,p,q)$-CompAWEC protocol $\Gamma$, we define $\Lambda(\kappa)={\Lambda^{\Gamma(\kappa)}(\kappa)}$. Assume towards a contradiction that $\Lambda$ is not a $(\alpha^{\prime},p^{\prime},q^{\prime})$-CompWEC. It follows that there exists a ppt $\mathsf{D}$ that for infinity many $\kappa\in{\mathbb{N}}$ contradicts one of the WEC secrecy properties of channel ensemble $\mathopen{}\mathclose{{}\left\{C_{\Lambda(\kappa)}}\right\}_{\kappa\in{\mathbb{N}}}$. Fix $\kappa\in{\mathbb{N}}$ for which this holds. By Lemma 4.4.4, there exists a ppt $\mathsf{E}^{\mathsf{D}}$ that for every such $\kappa$ contradicts one of the secrecy properties of the channel $C_{\Gamma(\kappa)}$. This implies that for infinity many $\kappa\in{\mathbb{N}}$, $\mathsf{E}^{\mathsf{D}}$ contradict the secrecy of the channel ensemble $\mathopen{}\mathclose{{}\left\{C_{\Gamma(\kappa)}}\right\}_{\kappa\in{\mathbb{N}}}$, which is a contradiction since this would means that $\Gamma$ is not a $(\ell,\alpha,p,q)$-CompAWEC.     $\Box$
4.5 Oblivious Transfer (OT)
4.5.0.1 Secure Computation.
We use the standard notion of securely computing a functionality, cf., [Goldreich04].
[…]
A two-party protocol securely computes a functionality $f$, if it does so according to the real/ideal paradigm. We add the term perfectly/statistically/computationally/non-uniform computationally, if the simulator’s output is perfect/statistical/computationally indistinguishable/ non-uniformly indistinguishable from the real distribution. The protocol have the above notions of security against semi-honest adversaries, if its security only guaranteed to holds against an adversary that follows the prescribed protocol. Finally, for the case of perfectly secure computation, we naturally apply the above notion also to the non-asymptotic case: the protocol with no security parameter perfectly compute a functionality $f$.
A two-party protocol securely computes a functionality ensemble $f$ with oracle to a channel $C$, if it does so according to the above definition when the parties have access to a trusted party computing $C$. All the above adjectives naturally extend to this setting.
4.5.0.2 Oblivious Transfer.
The (one-out-of-two) oblivious transfer functionality is defined as follows.
[…]
The oblivious transfer functionality over $\mathopen{}\mathclose{{}\left\{0,1}\right\}\times({\mathopen{}\mathclose{{}\left\{0,1}\right\}^{\ast}})^{2}$ is defined by $f_{{\sf OT}}(i,(\sigma_{0},\sigma_{1}))=(\perp,\sigma_{i})$.
A protocol is $\ast$ secure OT, for
[…]
We make use of the following useful results by Wullschleger on oblivious transfer amplification from weak channels.
Theorem 4.1 ([Wullschleger09], from WEC to statistically secure OT).
There exists an oracle aided protocol $\Pi$ such that the following holds: Given a $(\alpha,p,q)$-WEC $C$, if $44(\alpha+p)\leq 1-q$ then $\Pi^{C}(1^{\kappa})$ is a semi-honest statistically secure OT.
The following computational version of Theorem 4.1 is implicit in [Wullschleger09] and is based on the computational proof explicitly stated in [Wul07] (see Section 6 in [Wullschleger09] for discussion).
Theorem 4.2 ([Wullschleger09, Wul07], from CompWEC to computinally secure OT).
There exists an oracle aided protocol $\Pi$ such that the following holds: Given a $(\alpha,p,q)$-CompWEC protocol $\Lambda$, if $44(\alpha+p)\leq 1-q$ then $\Pi^{\Lambda}$ is a semi-honest computational secure OT.
5 Oblivious Transfer from ${\sf DP}$ Inner Product
In this section, we state our main theorems. We first state and prove the information theoretic case in Section 5.1, and then prove the computational case in Section 5.2.
5.1 Information-Theoretic Case
The following is our main theorem (for the information theoretic case) which shows that given a sufficiently accurate and private ${\sf DP}$ channel, we can construct a statistically secure semi-honest oblivious transfer protocol.
Theorem 5.1.
There exist constants $c_{1},c_{2}>0$ and an oracle-aided ppt protocol $\Pi$ such that the following holds for large enough $n\in{\mathbb{N}}$ and for $\varepsilon\leq\log^{0.9}n$, $\delta\leq\frac{1}{3n}$, and $\ell\leq e^{-c_{1}\varepsilon}c_{2}\cdot n^{1/6}$: Let $C=((X,U),(Y,V))$ be an $(\varepsilon,\delta)$-DP channel with independent $X,Y\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ that is $(\ell,0.999)$-accurate for the inner-product functionality (i.e., ${\mathrm{Pr}}_{C}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|O-\langle X,Y\rangle}\right|\leq\ell}\right]\geq 0.999$). Then $\Pi^{C}$ is a semi-honest statistically secure ${\sf OT}$ protocol.
To prove Theorem 5.1, we will make use of the following technical Lemma.
[…]
  1. 1.
    There exists an oracle-aided ppt algorithm ${\tilde{\mathsf{A}}}$ such that for every channel $C=((X,U),(Y,V))$ and algorithm $\mathsf{A}$ violating the AWEC secrecy property 1 of $\tilde{C}$ (the channel of $\Gamma^{C}$), the following holds for $Y^{*}_{i}={\tilde{\mathsf{A}}}^{\mathsf{A}}(i,\>Y_{-i},\>X,\>U)$:
    ${\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[Y^{*}_{i}=Y_{i}}\right]}\right]>e^{\varepsilon}\cdot{\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[Y^{*}_{i}=-Y_{i}}\right]}\right]+\delta.$
  2. 2.
    There exists an oracle-aided ppt algorithm ${\tilde{\mathsf{B}}}$ such that for every channel $C=((X,U),(Y,V))$ and algorithm $\mathsf{B}$ violating the AWEC secrecy property 2 of $\tilde{C}$ (the channel of $\Gamma^{C}$), the following holds for $X^{*}_{i}={\tilde{\mathsf{B}}}^{\mathsf{B},C}(i,\>X_{-i},\>Y,\>V)$:
    ${\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X^{*}_{i}=X_{i}}\right]}\right]>e^{\varepsilon}\cdot{\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X^{*}_{i}=-X_{i}}\right]}\right]+\delta.$
The proof of Section 5.1 is given in Section 6. But first we use Section 5.1 to prove Theorem 5.1.
Proof of Theorem 5.1.
By Section 5.1 there exists a ppt protocol $\Gamma\coloneqq\Gamma^{C}$ such that channel $\tilde{C}\coloneqq C_{\Gamma(\kappa)}$ is a $(\ell,\alpha=0.001,p=0.001,q=0.001)$-AWEC. By Section 4.4.4 (and Section 4.3) there exists a ppt protocol $\Lambda\coloneqq\Lambda^{\tilde{C}}$ such that the channel $\widehat{C}=C_{\Lambda(\kappa)}$ is a $(\alpha^{\prime}=0.002,\>p^{\prime}=0.001,\>q^{\prime}=1/2+0.022))$-WEC. Since $44(\alpha^{\prime}+p^{\prime})<1-q^{\prime}$, by Theorem 4.1 there exists a ppt protocol $\Pi$, such that $\Pi^{\widehat{C}}$ is a semi-honest statistically secure OT, concluding the proof.     $\Box$
5.2 Computational Case
In this section, we state and prove our results for the computational case. We show that a CDP (computational differential private) protocol that estimates the inner product well implies a semi-honest computationally secure oblivious transfer protocol.
Theorem 5.2 (Restatement of Theorem 1.1).
There exist constant $c_{1},c_{2}>0$ and an oracle-aided ppt protocol $\Pi$ such that the following holds for large enough $n\in{\mathbb{N}}$ and for $\varepsilon\leq\log^{0.9}n$, $\delta\leq\frac{1}{3n}$, and $\ell\leq e^{-c_{1}\varepsilon}c_{2}\cdot n^{1/6}$: Let $\Psi$ be an $(\varepsilon,\delta)$-CDP protocol that is $(\ell,0.999)$-accurate for the inner-product functionality. Then $\Pi^{\Psi}$ is a semi-honest computationally secure oblivious transfer protocol.
By the result of [GoldreichMW87], in the computational setting, we can “compile” any semi-honest computational oblivious transfer protocol into a protocol that is secure against any ppt (malicious) adversary (assuming one-way functions). We state this formally in the following corollary.
[…]
Let $\varepsilon,\delta,\ell$ be as in Theorem 5.2. If there exists a protocol $\Psi$ that is $(\varepsilon,\delta)$-CDP and is $(\ell,0.999)$-accurate for the inner-product functionality, then there exists a computationally secure oblivious transfer protocol.
Proof of Section 5.2.
By Theorem 5.2, there exists a semi-honest computational secure oblivious transfer protocol $\Pi$. Note that by [ImpagliazzoLu89], $\Pi$ implies the existence of one-way functions and by [GoldreichMW87], using the one-way function, we can compile $\Pi$ into a (computational) OT secure against arbitrary adversaries.     $\Box$
5.2.0.1 Proof of Theorem 5.2.
In order to use Section 5.1, and similar to [HaitnerMST22], we first convert the protocol $\Psi$ into a (no input) protocol such that the CDP-channel it induces, is uniform and accurately estimates the inner-product functionality. Such a transformation is simply the following protocol that invokes $\Psi$ over uniform inputs, and each party locally outputs its input.
[…]
Let $C$ be the channel ensemble induced by $\widehat{\Psi}$, letting its designated output (the function ${\rm out}$) be the designated output of the embedded execution of $\Psi$. The following fact is immediate by definition.
[…]
The channel ensemble $C$ is $(\varepsilon,\delta)$-${\sf CDP}$, and has the same accuracy for computing the inner product as protocol $\Psi$ has.
We first prove the computational version of Section 5.1.
[…]
There exist constant $c_{1},c_{2}>0$ and an oracle-aided ppt protocol $\Gamma=(\mathsf{A},\mathsf{B})$ such that the following holds for large enough $n\in{\mathbb{N}}$ and for $\varepsilon\leq\log^{0.9}n$, $\delta\leq\frac{1}{3n}$, and $\ell\leq e^{-c_{1}\varepsilon}c_{2}\cdot n^{1/6}$: Let $\Psi$ be an $(\varepsilon,\delta)$-CDP protocol that is $(\ell,0.999)$-accurate for the inner-product functionality. Then the protocol $\Gamma^{\Psi}$ is an $(\ell,\alpha=0.001,p=0.001,q=0.001)$-CompAWEC protocol.
Proof of 5.2.0.1.
Recall that $\Psi$ is an $(\varepsilon,\delta)$-${\sf DP}$ protocol that is $(\ell,0.999)$-acccurate for the inner-product functionality for $\varepsilon\leq\log^{0.9}n$, $\delta\leq\frac{1}{3n}$, and $\ell\leq e^{-c_{1}\varepsilon}c_{2}\cdot n^{1/6}$. By 5.2.0.1, $\widehat{\Psi}\coloneqq\widehat{\Psi}^{\Psi}$ induces a channel ensemble $C=\mathopen{}\mathclose{{}\left\{C_{\kappa}=((X_{\kappa},U_{\kappa}),(Y_{\kappa},V_{\kappa}))}\right\}_{\kappa\in{\mathbb{N}}}$ that is $(\varepsilon,\delta)$-CDP and $(\ell,0.999)$-acccurate for the inner-product functionality. Let $\Gamma$ be the ppt protocol guarrented by Section 5.1, we now claim that $\Gamma^{\Psi}(\kappa)\coloneqq\Gamma^{C_{\kappa}}(\kappa)$ is an $(\ell,\alpha=0.001,p=0.001,q=0.001)$-CompAWEC protocol. Assume towards contradiction that this does not hold, then by Section 5.1, there exists a ppt $\mathsf{D}$, such that for infinity many $\kappa\in{\mathbb{N}}$, $\mathsf{D}$ contradicts one of the two secrecy properties of ${\sf CompAWEC}$. Fix such $\kappa\in{\mathbb{N}}$ and omit it from the notation when clear from the context, and without loss of generality, assume that $\mathsf{D}$ contradicts the first secrecy property of ${\sf CompAWEC}$ (the case where $D$ contradicts the second property is essentially identical). By the first item of Section 5.1, there exists a ppt algorithm $\tilde{\mathsf{A}}$ such that $Y^{*}_{i}={\tilde{\mathsf{A}}}^{\mathsf{D}}(i,\>Y_{-i},\>X,\>U)$ and it holds that:
${\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[Y^{*}_{i}=Y_{i}}\right]}\right]>e^{\varepsilon}\cdot{\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[Y^{*}_{i}=-Y_{i}}\right]}\right]+\delta.$
Thus, we get a contradiction since by Section 4.3, algorithm ${\tilde{\mathsf{A}}}^{\mathsf{D}}$ breaks the CDP property of $C$.     $\Box$
Proof of Theorem 5.2.
Recall that $\Psi$ is an $(\varepsilon,\delta)$-${\sf DP}$ protocol that is $(\ell,0.999)$-acccurate for the inner-product functionality. By 5.2.0.1 there exists a ppt protocol $\Gamma$ such that $\Gamma\coloneqq\Gamma^{\Psi}$ is an $(\ell,\alpha=0.001,p=0.001,q=0.001)$-CompAWEC protocol. By Section 4.4.4, there exists a ppt protocol $\Lambda$ such that $\Lambda\coloneqq\Lambda^{\Gamma}$ is $(\epsilon^{\prime}=\epsilon+0.001,\>p^{\prime}=p,\>q^{\prime}=1/2+2(q+0.01))$-CompWEC. Finally, since $44(\varepsilon^{\prime}+p^{\prime})<1-q^{\prime}$ by Theorem 4.2 there exists a ppt protocol $\Pi$, such that $\Pi^{\Lambda}$ is a semi-honest computationally secure OT, as required.     $\Box$
6 AWEC From DP Inner Product (Proof of Section 5.1)
In this section, we show how to implement AWEC (Section 4.4.4) from an $(\varepsilon,\delta)$-DP channel that is accurate enough for the inner product functionality. We do that using a ppt protocol and a constructive security proof.
The following protocol is used to prove Section 5.1.
[…]
In the following, let $\Pi$ be Section 6, let $C=((X,U),(Y,V))$ be a uniform channel that is $(\ell,0.999)$-accurate for the inner-product functionality, let $\tilde{C}=((O_{A},V_{A}),(O_{B},V_{B}))$ be the channel that is induced by $\Pi^{C}$ (i.e., the parties’ outputs and views in the execution of $\Pi$ with oracle access to $C$), and let $R$, $\widetilde{Y}$, $\widehat{Y}$, $I_{1},\ldots,I_{k}$ be the random variables of the values of $r,\widetilde{y},\widehat{y},i_{1},\ldots,i_{k}$ in the execution of $\Pi^{C}$ (recall that the view of $\mathsf{B}$ in the execution is $V_{B}=(Y,V,R,X_{R},I_{1},\ldots,I_{k},\widetilde{Y})$, and the view of $\mathsf{A}$ is $V_{A}=(X,U,R,\widehat{Y}_{-R})$).
Recall that to prove Section 5.1, we need to prove that the channel $\tilde{C}$ satisfies the accuracy and secrecy properties of AWEC (Section 4.4.4), and in addition, the secrecy guarantees are constructive as stated in Properties 1-2 of Lemma 5.1 (i.e., a violation of at least one secrecy guaranty results with an efficient privacy attack on the DP channel $C$).
We first start with the easy part, which is to prove the accuracy guarantee of $\tilde{C}$.
[…]
It holds that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|O_{A}-O_{B}}\right|>\ell\mid O_{B}\neq\bot}\right]<0.001.$
Proof.
Compute
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|O_{A}-O_{B}}\right|>\ell\mid O_{B}\neq\bot}\right]$ $={\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|\langle X_{-R},Y_{-R}\rangle-\mathopen{}\mathclose{{}\left({\rm out}(V)-\langle X_{R},Y_{R}\rangle}\right)}\right|>\ell}\right]$
$={\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|{\rm out}(V)-\langle X,Y\rangle}\right|>\ell}\right]$
$<0.001,$
as required. The first equality holds since conditioned on $O_{B}\neq\bot$ it holds that $O_{A}=\langle X_{-R},Y_{-R}\rangle$, and the inequality holds since $C$ is an $(\ell,0.999)$-accurate for the inner-product functionality.     $\Box$
We next move to prove the secrecy guarantees of $\tilde{C}$ in a constructive manner. In Section 6.1 we prove property 1, and in Section 6.2 we prove property 2.
[…]
Finally, the following claim (proven in LABEL:sec:property_2) captures the second secrecy guarantee where party $\mathsf{B}$ cannot estimate $O_{A}$ too well, as otherwise, such an estimator $\mathsf{B}$ can be used to construct an efficient attack ${\tilde{\mathsf{B}}}$ that violates the privacy guarantee of the DP channel $C$.
[…]
There exists an oracle-aided ppt algorithm ${\tilde{\mathsf{B}}}$ such that for every algorithm $\mathsf{B}$ violating the AWEC secrecy property 2 of $\tilde{C}$, i.e.
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|\mathsf{B}(V_{B})-O_{A}}\right|\leq 1000\ell\mid O_{B}=\bot}\right]>q,$
where $\frac{131072\ell_{2}^{2}}{q^{2}k}\leq\frac{1}{e^{\varepsilon}+1}$, the for $X_{i}^{*}={\tilde{\mathsf{B}}}^{\mathsf{B}}(i,\>X_{-i},\>Y,\>V)$ it holds that
${\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X_{i}^{*}=X_{i}\mid X_{i}^{*}\neq\bot}\right]}\right]>\frac{e^{\varepsilon}}{e^{\varepsilon}+1}.$
Proof.
We remind that $\tilde{C}$ is an AWEC channel constructed by a differentially private inner-product protocol. Let =$\mathcal{R}=\{i|r_{i}=0\}$ and =${\cal I}=\mathopen{}\mathclose{{}\left\{i_{1},\dots,i_{k}}\right\}\cap{\cal R}$remove=$\mathcal{I}=\{i|i\in\{i_{1},\dots,i_{k}\},i\in\mathcal{R}\}$. Then $O_{A}=o-\langle x_{\mathcal{R}\setminus\mathcal{I}},\widehat{y}_{\mathcal{R}\setminus\mathcal{I}}\rangle-\langle x_{\mathcal{I}},\widehat{y}_{\mathcal{I}}\rangle$. It suffices to prove party $\mathsf{B}$ cannot predict $\langle x_{\mathcal{I}},\widehat{y}_{\mathcal{I}}\rangle$ with high accuracy. We put the details in Section LABEL:sec:property_2     $\Box$
6.1 B’s Security: Proving Property 1 of Section 5.1
Let $\mathsf{A}$ be an algorithm that violates the AWEC secrecy property 1 of $\tilde{C}$ (the channel of $\Pi^{C}$). Namely,
$\mathopen{}\mathclose{{}\left|{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{A}(V_{A})=1\mid O_{B}\neq\perp}\right]-{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{A}(V_{A})=1\mid O_{B}=\perp}\right]}\right|>\frac{1}{1000}.$
Recall that $V_{A}=(X,U,R,\widehat{Y}_{-R})$ where $\widehat{Y}=\begin{cases}Y&O_{B}\neq\perp\\ \widetilde{Y}&O_{B}=\perp\end{cases}$ and that $\widetilde{Y}$ is obtained from $Y$ by planting uniformly random $\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$ values in the random locations $I_{1},\ldots,I_{k}$ of $Y$ (the value of $k$ will be determined later by the analysis).
Therefore, the above inequality is equivalent to
$\mathopen{}\mathclose{{}\left|{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{A}(X,U,R,Y_{-R})=1}\right]-{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{A}(X,U,R,\widetilde{Y}_{-R})=1}\right]}\right|>\frac{1}{1000}.$
We assume without loss of generality that $\mathsf{A}$ outputs a $\mathopen{}\mathclose{{}\left\{0,1}\right\}$ bit, so the above inequality can be written as
$\mathopen{}\mathclose{{}\left|{\mathrm{E}}\mathopen{}\mathclose{{}\left[\mathsf{A}(X,U,R,Y_{-R})-\mathsf{A}(X,U,R,\widetilde{Y}_{-R})}\right]}\right|>\frac{1}{1000}.$ (10)
In the following, for $j\in\mathopen{}\mathclose{{}\left\{0,\ldots,k}\right\}$, let $\widetilde{Y}^{j}=(\widetilde{Y}^{j}_{1},\ldots,\widetilde{Y}^{j}_{n})$ where $\widetilde{Y}^{j}_{i}=\begin{cases}\widetilde{Y}_{i}&i\in\mathopen{}\mathclose{{}\left\{I_{1},\ldots,I_{j}}\right\}\\ Y_{i}&i\notin\mathopen{}\mathclose{{}\left\{I_{1},\ldots,I_{j}}\right\}\end{cases}$. Note that $\widetilde{Y}^{0}=Y$ and $\widetilde{Y}^{k}=\widetilde{Y}$. Therefore,
$\mathopen{}\mathclose{{}\left|{\mathrm{E}}_{j\leftarrow[k]}\mathopen{}\mathclose{{}\left[{\mathrm{E}}\mathopen{}\mathclose{{}\left[\mathsf{A}(X,U,R,Y^{(j-1)}_{-R})-\mathsf{A}(X,U,R,\widetilde{Y}^{(j)}_{-R})}\right]}\right]}\right|$ (11)
$=\frac{1}{k}\cdot\mathopen{}\mathclose{{}\left|\sum_{j=1}^{k}{\mathrm{E}}\mathopen{}\mathclose{{}\left[\mathsf{A}(X,U,R,Y^{(j-1)}_{-R})}\right]-{\mathrm{E}}\mathopen{}\mathclose{{}\left[\mathsf{A}(X,U,R,Y^{(j)}_{-R})}\right]}\right|$
$=\frac{1}{k}\cdot\mathopen{}\mathclose{{}\left|{\mathrm{E}}\mathopen{}\mathclose{{}\left[\mathsf{A}(X,U,R,Y_{-R})-\mathsf{A}(X,U,R,\widetilde{Y}_{-R})}\right]}\right|$
$>\frac{1}{1000k},$
where the inequality holds by Equation 10.
In the following, let $J\leftarrow[n]$ (sampled independently of the other random variables defined above), let $Z=(Z_{1},\ldots,Z_{n})=\widetilde{Y}^{J-1}$ and let $Z^{(i)}=(Z_{1},\ldots,Z_{i-1},-Z_{i},Z_{i+1},\ldots,Z_{n})$. It holds that
$\mathopen{}\mathclose{{}\left|{\mathrm{E}}\mathopen{}\mathclose{{}\left[\mathsf{A}(X,U,R,Z_{-R})}\right]-{\mathrm{E}}\mathopen{}\mathclose{{}\left[\mathsf{A}(X,U,R,Z^{(I_{J})}_{-R})}\right]}\right|$ (12)
$=\frac{1}{2}\cdot\mathopen{}\mathclose{{}\left|{\mathrm{E}}\mathopen{}\mathclose{{}\left[\mathsf{A}(X,U,R,Y^{(J-1)}_{-R})-\mathsf{A}(X,U,R,\widetilde{Y}^{(J)}_{-R})}\right]}\right|$
$=\frac{1}{2}\cdot\mathopen{}\mathclose{{}\left|{\mathrm{E}}_{j\leftarrow[k]}\mathopen{}\mathclose{{}\left[{\mathrm{E}}\mathopen{}\mathclose{{}\left[\mathsf{A}(X,U,R,Y^{(j-1)}_{-R})-\mathsf{A}(X,U,R,\widetilde{Y}^{(j)}_{-R})}\right]}\right]}\right|$
$>\frac{1}{2000k},$
where the first equality holds since $\widetilde{Y}^{J}$ is equal to $Z$ w.p. $1/2$ (happens when $\widetilde{Y}_{I_{J}}=Y_{I_{J}}$) and otherwise is equal to $Z^{(I_{J})}$ (happens when $\widetilde{Y}_{I_{J}}\neq Y_{I_{J}}$), and the last inequality holds by Equation 11.
We next use the following lemma, proven in Section C.2 (recall that a proof overview appears in Section 2.3).
[…]
Let $(Z,W)\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{0,1}\right\}^{*}$ be jointly distributed random variables, let $R\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}$ and $I\leftarrow[n]$ (sampled independently), and let $F$ be a (randomized) function that satisfies
$\mathopen{}\mathclose{{}\left|{\mathrm{E}}\mathopen{}\mathclose{{}\left[F(R,Z_{R},W)-F(R,Z^{(I)}_{R},W)}\right]}\right|\geq\gamma,$
for $Z^{(I)}=(Z_{1},\ldots,Z_{I-1},-Z_{I},Z_{I+1},\ldots,Z_{n})$. Then
  1. 1.
    ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}^{F}(I,Z_{-I},W)=-Z_{I}}\right]\leq O\mathopen{}\mathclose{{}\left(\frac{1}{\gamma^{2}n}}\right)$, and
  2. 2.
    ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}^{F}(I,Z_{-I},W)=Z_{I}}\right]\geq\Omega(\gamma)-O\mathopen{}\mathclose{{}\left(\frac{1}{\gamma^{2}n}}\right)$.
We now ready to finalize the proof of Property 1 of Section 5.1 using Equations 12 and 6.1.
Proof.
Consider the following oracle-aided algorithm ${\tilde{\mathsf{A}}}$:
[…]
Inputs: $i\in[n]$, $y_{-i}\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n-1}$, $x\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ and $u\in\mathopen{}\mathclose{{}\left\{0,1}\right\}^{*}$. Oracle: Deterministic algorithm $\mathsf{A}$. Operation: 1. Sample $j\leftarrow[k]$ and $i_{1},\ldots,i_{j-1}\leftarrow[n]$. 2. Sample $z_{-i}=(z_{1},\ldots,z_{i-1},z_{i+1},\ldots,z_{n})$, where for $t\in[n]\setminus\mathopen{}\mathclose{{}\left\{i}\right\}$: z_t ←{U{1,1}t ∈{i1,…,ij-1}ytotherwise. 3. Output $y_{i}^{*}\leftarrow\mathsf{G}^{F}(i,z_{-i},w)$ for $w=(x,u)$, where $\mathsf{G}=\mathsf{G}_{\frac{1}{2000k}}$ is the algorithm from Section 6.1, and $F(r,z_{r},w):=\mathsf{A}(w,(1-r_{1},\ldots,1-r_{n}),z_{r})$.
Note that by Equation 12 it holds that
$\mathopen{}\mathclose{{}\left|{\mathrm{E}}\mathopen{}\mathclose{{}\left[F(R,Z_{R},W)}\right]-{\mathrm{E}}\mathopen{}\mathclose{{}\left[F(R,Z_{R}^{I},W)}\right]}\right|>\frac{1}{2000k}$
for $W=(X,U)$, $I=I_{J}$ and the function $F(r,z_{r},w=(x,u)):=\mathsf{A}(w,(1-r_{1},\ldots,1-r_{n}),z_{r})$. Thus, Section 6.1 implies that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}^{F}(I,Z_{-I},V)=-Z_{I}}\right]\leq O\mathopen{}\mathclose{{}\left(\frac{k^{2}}{n}}\right)$ (13)
and
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}^{F}(I,Z_{-I},V)=Z_{I}}\right]\geq\Omega(1/k)-O\mathopen{}\mathclose{{}\left(\frac{k^{2}}{n}}\right).$ (14)
Recall that we denote $Y_{i}^{*}={\tilde{\mathsf{A}}}^{\mathsf{A}}(i,Y_{-i},X,U)$. We next lower-bound ${\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[Y_{i}^{*}=Y_{i}}\right]}\right]$ and upper-bound ${\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[Y_{i}^{*}=-Y_{i}}\right]}\right]$. The first bound hold by the following calculation
${\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[Y_{i}^{*}=Y_{i}}\right]}\right]$ $\geq 0.9\cdot{\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[Y_{i}^{*}=Z_{i}}\right]}\right]$ (15)
$=0.9\cdot{\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}^{F}(i,Z_{-i},W)=Z_{i}}\right]}\right]$
$\geq\Omega\mathopen{}\mathclose{{}\left(\frac{1}{k}}\right)-O\mathopen{}\mathclose{{}\left(\frac{k^{2}}{n}}\right).$
The first inequality holds since $i\notin\mathopen{}\mathclose{{}\left\{I_{1},\ldots,I_{J-1}}\right\}\implies Z_{i}=Y_{i}$ and ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[i\notin\mathopen{}\mathclose{{}\left\{I_{1},\ldots,I_{J-1}}\right\}}\right]\geq\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{k-1}\geq 0.9$ (recall that $k\in o(n)$). The equality holds since, conditioned on $X=x,U=u,Y_{-i}=y_{-i}$ (the inputs of ${\tilde{\mathsf{A}}}$), the value of $z_{-i}$ that is sampled in Step 2 of ${\tilde{\mathsf{A}}}$ is distributed the same as $Z_{-i}|_{X=x,U=u,Y_{-i}=y_{-i}}$, and therefore, $Y_{i}^{*}\equiv\mathsf{G}^{F}(i,Z_{-i},W)$ for $W=(X,U)$. The last inequality holds by Equation 14.
On the other hand, we have that
${\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[Y_{i}^{*}=-Y_{i}}\right]}\right]$ $\leq{\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[Y_{i}^{*}=-Y_{i}\mid Y_{i}=Z_{i}}\right]+{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[Z_{i}\neq Y_{i}}\right]}\right]$ (16)
$\leq{\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[Y_{i}^{*}=-Z_{i}}\right]}\right]+\frac{k}{2n}$
$={\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}^{F}(i,Z_{-i})=-Z_{i}}\right]}\right]+\frac{k}{2n}$
$\leq O\mathopen{}\mathclose{{}\left(\frac{k^{2}}{n}}\right).$
The second inequality holds since for any fixing $Y_{-i}=y_{-i},X=x,U=u$, we have $Y_{i}^{*}={\tilde{\mathsf{A}}}^{\mathsf{A}}(i,y_{-i},x,u)$ which is independent of $(Y_{i},Z_{i})$, and also since $i\notin\mathopen{}\mathclose{{}\left\{I_{1},\ldots,I_{J-1}}\right\}\implies Z_{i}=Y_{i}$ which yields that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[Z_{i}\neq Y_{i}}\right]\leq 1-{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[i\notin\mathopen{}\mathclose{{}\left\{I_{1},\ldots,I_{J-1}}\right\}}\right]\leq 1-\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)^{k}\leq 1-e^{\frac{k}{2n}}\leq\frac{k}{2n}.$
The equality in Equation 16 holds since, conditioned on $X=x,U=u,Y_{-i}=y_{-i}$ (the inputs of ${\tilde{\mathsf{A}}}$), the value of $z_{-i}$ that is sampled in Step 2 of ${\tilde{\mathsf{A}}}$ is distributed the same as $Z_{-i}|_{X=x,U=u,Y_{-i}=y_{-i}}$, and therefore, $Y_{i}^{*}\equiv\mathsf{G}^{F}(i,Z_{-i},W)$ for $W=(X,U)$. The last inequality in Equation 16 holds by Equation 13.
By Equations 15 and 16, assuming that $\delta\leq 1/n$ where $n$ is large enough, there exists a constant $c>0$ such that if $k\leq c\cdot(e^{-\varepsilon}n)^{1/3}$ then ${\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[Y_{i}^{*}=Y_{i}}\right]}\right]>e^{\varepsilon}\cdot{\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[Y_{i}^{*}\neq Y_{i}}\right]}\right]+\delta$, as required. Recall that $k=\mathopen{}\mathclose{{}\left\lceil e^{\lambda_{1}\varepsilon}\lambda_{2}\cdot\ell^{2}}\right\rceil$, where $\lambda_{1},\lambda_{2}$ are the constants from Equation 28. Hence, we can set a bound of $e^{-c_{1}\varepsilon}c_{2}\cdot n^{1/6}$ on $\ell$ for $c_{1}=\lambda_{1}/2+1/6$ and $c_{2}=\sqrt{c/\lambda_{2}}$ to guarantee that $k\leq c\cdot(e^{-\varepsilon}n)^{1/3}$.
$\Box$
[…]
6.1.1 Proving Section 6.1
Proof.
Let $0<\gamma\leq 0.01$ and consider the following algorithm $\mathsf{G}=\mathsf{G}_{\gamma}$:
[…]
Inputs: $i\in[n]$, $z_{-i}=(z_{1},\ldots,z_{i-1},z_{i+1},\ldots,z_{n})\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n-1}$ and $w\in\mathopen{}\mathclose{{}\left\{0,1}\right\}^{*}$. Parameter: $0<\gamma\leq 0.01$. Oracle: $f\colon\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{\leq n}\times\mathopen{}\mathclose{{}\left\{0,1}\right\}^{*}\rightarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}$. Operation: 1. For $b\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$: (a) Let $z^{b}=(z_{1},\ldots,z_{i-1},b,z_{i+1},\ldots,z_{n})$. (b) Estimate $\mu^{b}:={\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\mid r_{i}=1,\>f}\mathopen{}\mathclose{{}\left[f(r,z^{b}_{r},w)}\right]$ as follows: Sample $r_{1},\ldots,r_{s}\leftarrow\mathopen{}\mathclose{{}\left\{r\in\mathopen{}\mathclose{{}\left\{0,1}\right\}\colon r_{i}=1}\right\}$, for $s=\mathopen{}\mathclose{{}\left\lceil\frac{128\log(6n)}{\gamma^{2}}}\right\rceil$, and then sample $\tilde{\mu}^{b}\leftarrow\frac{1}{s}\sum_{j=1}^{s}f(r_{j},z^{b}_{r_{j}},w)$ (using $s$ calls to the oracle $f$). 2. Estimate $\mu^{*}:={\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\mid r_{i}=0,\>f}\mathopen{}\mathclose{{}\left[f(r,z^{1}_{r},w)}\right]$ as follows: Sample $r_{1},\ldots,r_{s}\leftarrow\mathopen{}\mathclose{{}\left\{r\in\mathopen{}\mathclose{{}\left\{0,1}\right\}\colon r_{i}=0}\right\}$, for $s=\mathopen{}\mathclose{{}\left\lceil\frac{128\log(6n)}{\gamma^{2}}}\right\rceil$, and then sample $\tilde{\mu}^{*}\leftarrow\frac{1}{s}\sum_{j=1}^{s}f(r_{j},z^{1}_{r_{j}},w)$ (using $s$ calls to the oracle $f$). 3. If exists $b\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$ s.t. $\mathopen{}\mathclose{{}\left|\tilde{\mu}^{b}-\tilde{\mu}^{*}}\right|<\gamma/4$ and $\mathopen{}\mathclose{{}\left|\tilde{\mu}^{-b}-\tilde{\mu}^{*}}\right|>\gamma/4$, output $b$. 4. Otherwise, output $\bot$.
In the following, fix a pair of (jointly distributed) random variables $(Z,W)\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{0,1}\right\}^{*}$ and a randomized function $f\colon\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{\leq n}\times\mathopen{}\mathclose{{}\left\{0,1}\right\}^{*}\rightarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$ that satisfy
$\mathopen{}\mathclose{{}\left|{\mathrm{E}}\mathopen{}\mathclose{{}\left[f(R,Z_{R},W)-f(R,Z^{(I)}_{R},W)}\right]}\right|\geq\gamma,$
for $R\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}$ and $I\leftarrow[n]$ that are sampled independently. Our goal is to prove that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}^{f}(I,Z_{-I},W)=-Z_{I}}\right]\leq O\mathopen{}\mathclose{{}\left(\frac{1}{\gamma^{2}n}}\right),$ (17)
and
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}^{f}(I,Z_{-I},W)=Z_{I}}\right]\geq\Omega(\gamma)-O\mathopen{}\mathclose{{}\left(\frac{1}{\gamma^{2}n}}\right).$ (18)
Note that
$\text{For every random variable }D\in[-1,1]\text{ with }\mathopen{}\mathclose{{}\left|{\mathrm{E}}\mathopen{}\mathclose{{}\left[D}\right]}\right|\geq\gamma>0:\quad{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|D}\right|>\gamma/2}\right]>\gamma/2,$ (19)
as otherwise, $\mathopen{}\mathclose{{}\left|{\mathrm{E}}\mathopen{}\mathclose{{}\left[D}\right]}\right|\leq{\mathrm{E}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|D}\right|}\right]\leq 1\cdot\frac{\gamma}{2}+\frac{\gamma}{2}(1-\frac{\gamma}{2})<\gamma$.
By applying Equation 36 with $D={\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n},f}\mathopen{}\mathclose{{}\left[f(r,Z_{r},W)-f(r,Z^{(I)}_{r},W)}\right]$, it holds that
${\mathrm{Pr}}_{(i,z,w)\leftarrow(I,Z,W)}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|{\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n},f}\mathopen{}\mathclose{{}\left[f(r,z_{r},w)-f(r,z^{(i)}_{r},w)}\right]}\right|>\gamma/2}\right]>\gamma/2.$ (20)
On the other hand, for every fixing of $(z,w)\in\operatorname{Supp}(Z,W)$, we can apply Section C.2 with the function $f_{z,w}(r)=f(r,z_{r},w)$ and with $\alpha=\frac{\gamma}{16}$ to obtain that
${\mathrm{Pr}}_{i\leftarrow I}\mathopen{}\mathclose{{}\left[\>\mathopen{}\mathclose{{}\left|{\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\mid r_{i}=0,\>f}\mathopen{}\mathclose{{}\left[f(r,z_{r},w)}\right]-{\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\mid r_{i}=1,\>f}\mathopen{}\mathclose{{}\left[f(r,z_{r},w)}\right]\>}\right|\geq\frac{\gamma}{16}}\right]\leq\frac{512}{n\gamma^{2}}.$
But since the above holds for every fixing of $(z,w)$, then in particular it holds that
${\mathrm{Pr}}_{(i,z,w)\leftarrow(I,Z,W)}\mathopen{}\mathclose{{}\left[\>\mathopen{}\mathclose{{}\left|{\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\mid r_{i}=0,\>f}\mathopen{}\mathclose{{}\left[f(r,z_{r},w)}\right]-{\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\mid r_{i}=1,\>f}\mathopen{}\mathclose{{}\left[f(r,z_{r},w)}\right]\>}\right|\geq\frac{\gamma}{16}}\right]\leq\frac{512}{n\gamma^{2}}.$ (21)
We next prove Equations 34 and 35 using Equations 37 and 38.
In the following, for a triplet $t=(i,z,w)\in\operatorname{Supp}(I,Z,W)$, consider a random execution of $\mathsf{G}^{f}(i,z_{-i},w)$. For $x\in\mathopen{}\mathclose{{}\left\{-1,1,*}\right\}$, let $\mu^{x}_{t}$ be the value of $\mu^{x}$ in the execution, and let $M^{x}_{t}$ be the (random variable of the) value of $\tilde{\mu}^{x}$ in the execution. Note that by definition, it holds that $\mu_{i,z,w}^{z_{i}}={\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n},f}\mathopen{}\mathclose{{}\left[f(r,z_{r},w)}\right]$ and $\mu_{i,z,w}^{-z_{i}}={\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n},f}\mathopen{}\mathclose{{}\left[f(r,z^{(i)}_{r},w)}\right]$. Therefore, Equation 37 is equivalent to
${\mathrm{Pr}}_{t=(i,z,w)\leftarrow(I,Z,W)}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|\mu_{t}^{z_{i}}-\mu_{t}^{-z_{i}}}\right|>\gamma/2}\right]>\gamma/2.$ (22)
Furthermore, note that $\mu_{i,z,w}^{*}:={\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\mid r_{i}=0,\>f}\mathopen{}\mathclose{{}\left[f(r,z^{1}_{r},w)}\right]={\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\mid r_{i}=0,\>f}\mathopen{}\mathclose{{}\left[f(r,z_{r},w)}\right]$ and that $\mu_{i,z,w}^{z_{i}}={\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\mid r_{i}=1,\>f}\mathopen{}\mathclose{{}\left[f(r,z^{z_{i}}_{r},w)}\right]$. Therefore, Equation 38 is equivalent to
${\mathrm{Pr}}_{t=(i,z,w)\leftarrow(I,Z,W)}\mathopen{}\mathclose{{}\left[\>\mathopen{}\mathclose{{}\left|\mu_{t}^{*}-\mu_{t}^{z_{i}}}\right|\geq\frac{\gamma}{16}}\right]\leq\frac{512}{n\gamma^{2}}.$ (23)
We next prove the lemma using Equations 39 and 40.
Note that by Hoeffding’s inequality, for every $t=(i,z,w)\in\operatorname{Supp}(I,Z,W)$ and $x\in\mathopen{}\mathclose{{}\left\{-1,1,*}\right\}$ it holds that ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|M_{t}^{x}-\mu_{t}^{x}}\right|\geq\frac{\gamma}{16}}\right]\leq 2\cdot e^{-2s\mathopen{}\mathclose{{}\left(\frac{\gamma}{16}}\right)^{2}}\leq\frac{1}{3n}$, which yields that for every fixing of $t=(i,z,w)\in\operatorname{Supp}(I,Z,W)$, w.p. at least $1-1/n$ we have for all $x\in\mathopen{}\mathclose{{}\left\{-1,1,*}\right\}$ that $\mathopen{}\mathclose{{}\left|M_{t}^{x}-\mu_{t}^{x}}\right|<\frac{\gamma}{16}$ (denote this event by $E_{t}$).
The proof of Equation 34 holds by the following calculation:
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}^{f}(I,Z_{-I},W)=-Z_{I}}\right]$
$={\mathrm{E}}_{(i,z,w)\leftarrow(I,Z,W)}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}^{f}(i,z_{-i},w)=-z_{i}}\right]}\right]$
$={\mathrm{E}}_{t=(i,z,w)\leftarrow(I,Z,W)}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left|M_{t}^{*}-M_{t}^{z_{i}}}\right|>\gamma/4}\right\}\land\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left|M_{t}^{*}-M_{t}^{-z_{i}}}\right|<\gamma/4}\right\}}\right]}\right]$
$\leq{\mathrm{E}}_{t=(i,z,w)\leftarrow(I,Z,W)}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|M_{t}^{*}-M_{t}^{z_{i}}}\right|>\gamma/4}\right]}\right]$
$\leq{\mathrm{E}}_{t=(i,z,w)\leftarrow(I,Z,W)}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|M_{t}^{*}-M_{t}^{z_{i}}}\right|>\gamma/4\mid E_{t}}\right]}\right]+\frac{1}{n}$
$\leq{\mathrm{Pr}}_{t=(i,z,w)\leftarrow(I,Z,W)}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|\mu_{t}^{*}-\mu_{t}^{z_{i}}}\right|\geq\frac{\gamma}{16}}\right]+\frac{1}{n}$
$\leq\frac{512}{n\gamma^{2}}+\frac{1}{n},$
The second inequality holds since ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\neg E_{t}}\right]\leq 1/n$ for every $t$. The penultimate inequality holds since conditioned on $E_{t}$, it holds that $\mathopen{}\mathclose{{}\left|M_{t}^{*}-\mu_{t}^{*}}\right|\leq\frac{\gamma}{16}$ and $\mathopen{}\mathclose{{}\left|M_{t}^{z_{i}}-\mu_{t}^{z_{i}}}\right|\leq\frac{\gamma}{16}$, which implies that $\mathopen{}\mathclose{{}\left|M_{t}^{*}-M_{t}^{z_{i}}}\right|>\gamma/4\>\implies\>\mathopen{}\mathclose{{}\left|\mu_{t}^{*}-\mu_{t}^{z_{i}}}\right|\geq\frac{\gamma}{4}-2\cdot\frac{\gamma}{16}>\frac{\gamma}{16}$. The last inequality holds by Equation 40.
It is left to prove Equation 35. Observe that Equations 39 and 40 imply with probability at least $\gamma/2-\frac{512}{n\gamma^{2}}$ over $t=(i,z,w)\in(I,Z,W)$, we have $\mathopen{}\mathclose{{}\left|\mu_{t}^{*}-\mu_{t}^{z_{i}}}\right|\leq\frac{\gamma}{16}$ and $\mathopen{}\mathclose{{}\left|\mu_{t}^{z_{i}}-\mu_{t}^{-z_{i}}}\right|\geq\frac{\gamma}{2}$. Therefore, conditioned on event $E_{t}$ (i.e., $\forall x\in\mathopen{}\mathclose{{}\left\{-1,1,*}\right\}:\>\mathopen{}\mathclose{{}\left|M_{t}^{x}-\mu_{t}^{x}}\right|\leq\frac{\gamma}{16}$), we have for such triplets $t=(i,z,w)$ that
$\mathopen{}\mathclose{{}\left|M_{t}^{*}-M_{t}^{-z_{i}}}\right|$ $\geq\mathopen{}\mathclose{{}\left|\mu_{t}^{z_{i}}-\mu_{t}^{-z_{i}}}\right|-\mathopen{}\mathclose{{}\left|\mu_{t}^{z_{i}}-\mu_{t}^{*}}\right|-\mathopen{}\mathclose{{}\left|M_{t}^{*}-\mu_{t}^{*}}\right|-\mathopen{}\mathclose{{}\left|M_{t}^{-z_{i}}-\mu_{t}^{-z_{i}}}\right|$
$\geq\frac{\gamma}{2}-3\cdot\frac{\gamma}{16}$
$>\gamma/4,$
and
$\mathopen{}\mathclose{{}\left|M_{t}^{*}-M_{t}^{z_{i}}}\right|$ $\leq\mathopen{}\mathclose{{}\left|M_{t}^{*}-\mu_{t}^{*}}\right|+\mathopen{}\mathclose{{}\left|\mu_{t}^{*}-\mu_{t}^{z_{i}}}\right|+\mathopen{}\mathclose{{}\left|\mu_{t}^{z_{i}}-M_{t}^{z_{i}}}\right|$
$\leq 3\cdot\frac{\gamma}{16}$
$<\gamma/4.$
Thus, we conclude that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}^{f}(I,Z_{-I},W)=Z_{I}}\right]$
$={\mathrm{E}}_{(i,z,w)\leftarrow(I,Z,W)}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}^{f}(i,z_{-i},w)=z_{i}}\right]}\right]$
$={\mathrm{E}}_{t=(i,z,w)\leftarrow(I,Z,W)}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left|M_{t}^{*}-M_{t}^{-z_{i}}}\right|>\gamma/4}\right\}\land\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left|M_{t}^{*}-M_{t}^{z_{i}}}\right|<\gamma/4}\right\}}\right]}\right]$
$\geq\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)\cdot{\mathrm{E}}_{t=(i,z,w)\leftarrow(I,Z,W)}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left|M_{t}^{*}-M_{t}^{-z_{i}}}\right|>\gamma/4}\right\}\land\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left|M_{t}^{*}-M_{t}^{z_{i}}}\right|<\gamma/4}\right\}\mid E_{t}}\right]}\right]$
$\geq\mathopen{}\mathclose{{}\left(1-\frac{1}{n}}\right)\cdot\mathopen{}\mathclose{{}\left(\gamma/2-\frac{512}{n\gamma^{2}}}\right)$
$\geq\gamma/4-\frac{1024}{n\gamma^{2}},$
which proves Equation 35. The first inequality holds since ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[E_{t}}\right]\geq 1-1/n$ for every $t=(i,z,w)$, and the second one holds by the observation above.
$\Box$
6.2 A’s Security: Proving Property 2 of Section 5.1
Let $\mathsf{B}$ be an algorithm that violates the AWEC secrecy property 2 of $\tilde{C}=((V_{A},O_{A}),(V_{B},O_{B}))$ — the channel of $\Pi^{C=((X,U),(Y,V))}$ (Section 6). Namely,
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|\mathsf{B}(V_{B})-O_{A}}\right|\leq 1000\ell\mid O_{B}=\bot}\right]>\frac{1}{1000}.$ (24)
Recall that $V_{B}=(Y,V,R,X_{R},I_{1},\ldots,I_{k},\widetilde{Y})$ where $I_{1},\ldots,I_{k}\leftarrow[n]$ are the indices that $\mathsf{B}$ chooses at Step 3b, and $\widetilde{Y}=(\widetilde{Y}_{1},\ldots,\widetilde{Y}_{n})$ where $\widetilde{Y}_{i}\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$ for $i\in\mathopen{}\mathclose{{}\left\{I_{1},\ldots,I_{k}}\right\}$ and otherwise $\widetilde{Y}_{i}=Y_{i}$. Furthermore, conditioned on $O_{B}=\bot$, recall that $O_{A}=\langle X_{-R},\widetilde{Y}_{-R}\rangle$. Therefore, Equation 24 is equivalent to
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|\mathsf{B}(V_{B})-\langle X_{-R},\widetilde{Y}_{-R}\rangle}\right|\leq 1000\ell}\right]>\frac{1}{1000}.$ (25)
In the following, define the random variable $H$ to be the first $m=\mathopen{}\mathclose{{}\left\lceil k/4}\right\rceil$ indices of $R^{0}\cap\mathopen{}\mathclose{{}\left\{I_{1},\ldots,I_{k}}\right\}$ for $R^{0}=\mathopen{}\mathclose{{}\left\{i\colon R_{i}=0}\right\}$, where we let $H=\emptyset$ if the size of the intersection is smaller than $m$. Since $R\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}$, Hoeffding’s inequality implies that ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|R^{0}}\right|\geq 0.4n}\right]\geq 1-e^{-\Omega(n)}$. Since $I_{1},\ldots,I_{k}\leftarrow[n]$ (independent of $R$), then again by Hoeffding’s inequality we obtain that ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|H}\right|=\mathopen{}\mathclose{{}\left\lceil k/4}\right\rceil\mid\mathopen{}\mathclose{{}\left|R^{0}}\right|\geq 0.4n}\right]\geq 1-e^{-\Omega(k)}$, which yields that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[H\neq\emptyset}\right]={\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|H}\right|=\mathopen{}\mathclose{{}\left\lceil k/4}\right\rceil}\right]\geq 1-e^{-\Omega(k)}-e^{-\Omega(n)}\geq 1-\frac{1}{10000}.$
Therefore, by the union bound,
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left|\mathsf{B}(V_{B})-\langle X_{-R},\widetilde{Y}_{-R}\rangle}\right|\leq 1000\ell}\right\}\land\mathopen{}\mathclose{{}\left\{H\neq\emptyset}\right\}}\right]$ (26)
$=1-{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left|\mathsf{B}(V_{B})-\langle X_{-R},\widetilde{Y}_{-R}\rangle}\right|>1000\ell}\right\}\lor\mathopen{}\mathclose{{}\left\{H=\emptyset}\right\}}\right]$
$\geq 1-{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left|\mathsf{B}(V_{B})-\langle X_{-R},\widetilde{Y}_{-R}\rangle}\right|>1000\ell}\right\}}\right]-{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\{H=\emptyset}\right\}}\right]$
$\geq 1-\mathopen{}\mathclose{{}\left(1-\frac{1}{1000}}\right)-\frac{1}{10000}$
$\geq\frac{1}{2000}.$
In the following, let $d$ be the number of random coins that $\mathsf{B}$ uses, and for $\psi\in\mathopen{}\mathclose{{}\left\{0,1}\right\}^{d}$ let $\mathsf{B}_{\psi}$ be algorithm $\mathsf{B}$ when fixing its random coins to $\psi$. Let $\Psi\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{d}$, $Z=X_{H}$, $T^{\prime}=(Y,V,R,I_{1},\ldots,I_{k},H,\widetilde{Y}_{-{\cal H}},X_{-{\cal H}})$, $T=(\Psi,T^{\prime})$ and $S=\widetilde{Y}_{H}$. Note that conditioned on $H\neq\perp$, $S$ is a uniformly random string in $\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{m}$, independent of $Z$ and $T$, and note that $V_{B}$ is a deterministic function of $(\widetilde{Y}_{H},T)$ (because $X_{R}$, which is part of $V_{B}$, is a sub-string of $X_{-{\cal H}}$). Equation 26 yields that w.p. $1/2000$ over $z\leftarrow Z$, $t=(\psi,t^{\prime}=(y,v,r,i_{1},\ldots,i_{k},{\cal H},x_{-{\cal H}},\widetilde{y}_{-{\cal H}}))\leftarrow T$ and $s\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{m}$, the following holds for ¯=$\bar{{\cal H}}=\mathopen{}\mathclose{{}\left\{i\in[n]\colon r_{i}=0}\right\}\setminus{\cal H}\>$:
$\mathopen{}\mathclose{{}\left|\mathsf{B}_{\psi}(s,t^{\prime})-\langle x_{\bar{{\cal H}}},\widetilde{y}_{\bar{{\cal H}}}\rangle-\langle z,s\rangle}\right|\leq 1000\ell.$
By denoting $f(s,t=(\psi,t^{\prime}))=\mathsf{B}_{\psi}(s,t^{\prime})+\langle x_{\bar{{\cal H}}},\widetilde{y}_{\bar{{\cal H}}}\rangle$, [^12] Note that $f(s,t)$ is well-defined because $t$ contains $x_{\bar{{\cal H}}}$ and $\widetilde{y}_{\bar{{\cal H}}}$ (sub-strings of $x_{-{\cal H}}$ and $\widetilde{y}_{-{\cal H}}$, respectively). the above observation is equivalent to
${\mathrm{Pr}}_{(z,t)\leftarrow(Z,T),\>s\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{m}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|f(s,t)-\langle z,s\rangle}\right|\leq 1000\ell}\right]\geq\frac{1}{2000}.$ (27)
In the following, let ${\cal D}$ be the joint distribution of $(Z,T)$, which is equivalent to the output of $\mathsf{GenView}^{C}()$ defined below in Section 6.2.
[…]
1. Sample $\psi\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{d}$. 2. Sample $r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}$ and $i_{1},\ldots,i_{k}\leftarrow[n]$. 3. Let ${\cal H}$ be the first $m=\mathopen{}\mathclose{{}\left\lceil k/4}\right\rceil$ indices of $\mathopen{}\mathclose{{}\left\{i\in[n]\colon r_{i}=0}\right\}\cap\mathopen{}\mathclose{{}\left\{i_{1},\ldots,i_{k}}\right\}$, where ${\cal H}=\emptyset$ if the intersection size is less than $m$. 4. Let ¯=$\bar{{\cal H}}=\mathopen{}\mathclose{{}\left\{i\in[n]\colon r_{i}=0}\right\}\setminus{\cal H}$. 5. Sample $\widetilde{y}_{\bar{{\cal H}}}\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{\mathopen{}\mathclose{{}\left|\bar{{\cal H}}}\right|}$. 6. Output $(\psi,r,i_{1},\ldots,i_{k},{\cal H},\widetilde{y}_{\bar{{\cal H}}})$.
[…]
Oracle: A channel $C=((X,U),(Y,V))$. Operation: 1. Sample $((x,u),(y,v))\leftarrow C$. 2. Sample $(\psi,r,i_{1},\ldots,i_{k},{\cal H},\widetilde{y}_{\bar{{\cal H}}})\leftarrow\mathsf{GenRand}()$ (Section 6.2). 3. Output $z=x_{{\cal H}}$ and $t=(\psi,t^{\prime})$ for $t^{\prime}=(y,v,r,i_{1},\ldots,i_{k},{\cal H},\widetilde{y}_{-{\cal H}},x_{-{\cal H}})$, where $\widetilde{y}_{i}=y_{i}$ for $i\in[n]\setminus({\cal H}\cup\bar{{\cal H}})$.
We now can use the following reconstruction result from [HaitnerMST22]:
[…]
There exists constants $\eta_{1},\eta_{2}>0$ and a ppt algorithm $\mathsf{Dist}$ such that the following holds for large enough $m\in{\mathbb{N}}$: Let $\varepsilon\geq 0$ and $a\geq\log m$, and let ${\cal D}$ be a distribution over $\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{m}\times\Sigma^{*}$. Then for every function $f$ that satisfies
${\mathrm{Pr}}_{(z,t)\leftarrow{\cal D},\>s\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{m}}\mathopen{}\mathclose{{}\left[f(s,t)-\langle z,s\rangle\leq a}\right]\geq e^{\eta_{1}\cdot\varepsilon}\cdot\eta_{2}\cdot a/\sqrt{m},$
it holds that
${\mathrm{Pr}}_{(z,t)\leftarrow{\cal D},\>j\leftarrow[m]}\mathopen{}\mathclose{{}\left[\mathsf{Dist}^{{\cal D},f}(j,z,t)=1}\right]>e^{\varepsilon}\cdot{\mathrm{Pr}}_{(z,t)\leftarrow{\cal D},\>j\leftarrow[m]}\mathopen{}\mathclose{{}\left[\mathsf{Dist}^{{\cal D},f}(j,z^{(j)},t)=1}\right]+\frac{1}{m},$
where $z^{(j)}=(z_{1},\ldots,z_{j-1},-z_{j},z_{j+1},\ldots,z_{m})$. [^13] Theorem 4.6 in [HaitnerMST22] actually considered a harder setting where $z$ is the coordinate-wise product of two vectors $x,y\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$, and $f$ only guarantees accuracy when in addition to $s$ and $t$, it also gets $x_{s}=\mathopen{}\mathclose{{}\left\{x_{i}\colon s_{i}=1}\right\}$ and $y_{-s}=\mathopen{}\mathclose{{}\left\{y_{i}\colon s_{i}=-1}\right\}$ as inputs.
Now, we would like to apply Section 6.2 with ${\cal D}=\mathsf{GenView}^{C}()$ and $a=1000\ell$. To do that, Section 6.2 yields that we need to choose $k$ such that $\frac{e^{\eta_{1}\cdot\varepsilon}\cdot\eta_{2}\cdot 1000\ell}{\mathopen{}\mathclose{{}\left\lceil k/4}\right\rceil}\leq\frac{1}{2000}$, which holds by choosing $k=\mathopen{}\mathclose{{}\left\lfloor e^{\lambda_{1}\varepsilon}\cdot\lambda_{2}\cdot\ell^{2}}\right\rfloor$ with
$\lambda_{1}=\eta_{1}\text{ and }\lambda_{2}=10^{7}\eta_{2}+1,$ (28)
where $\eta_{1},\eta_{2}$ are the constants from Section 6.2.
We deduce from Sections 6.2 and 27 that
${\mathrm{Pr}}_{(z,t)\leftarrow{\cal D},\>j\leftarrow[m]}\mathopen{}\mathclose{{}\left[\mathsf{Dist}^{{\cal D},f}(j,z,t)=1}\right]>e^{\varepsilon}\cdot{\mathrm{Pr}}_{(z,t)\leftarrow{\cal D},\>j\leftarrow[m]}\mathopen{}\mathclose{{}\left[\mathsf{Dist}^{{\cal D},f}(j,z^{(j)},t)=1}\right]+\frac{1}{m}.$ (29)
We now ready to describe our algorithm ${\tilde{\mathsf{B}}}$ that satisfies Property 2 of Section 5.1.
[…]
Inputs: $i\in[n]$, $x_{-i}=(x_{1},\ldots,x_{i-1},x_{i+1},\ldots,x_{n})\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n-1}$, $y\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ and $v\in\mathopen{}\mathclose{{}\left\{0,1}\right\}^{*}$. Oracles: A channel $C=((X,U),(Y,V))$ and an algorithm $\mathsf{B}$. Operation: 1. Sample $(\psi,r,i_{1},\ldots,i_{k},{\cal H},\widetilde{y}_{\bar{{\cal H}}})\leftarrow\mathsf{GenRand}()$ (Section 6.2). 2. If $i\notin{\cal H}$, output $\bot$. 3. Otherwise: (a) Let $t=(y,v,r,i_{1},\ldots,i_{k},{\cal H},\widetilde{y}_{-{\cal H}},x_{-{\cal H}})$ where $\widetilde{y}_{i^{\prime}}=y_{i^{\prime}}$ for $i^{\prime}\in[n]\setminus({\cal H}\cup\bar{{\cal H}})$. (b) For $b\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$: Let $x^{b}=(x_{1},\ldots,x_{i-1},b,x_{i+1},\ldots,x_{n})$ and $z^{b}=x^{b}_{{\cal H}}\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{m}$, where $m=\mathopen{}\mathclose{{}\left\lceil k/4}\right\rceil$. (c) Let $j\in[m]$ be the index such that $z^{1}_{j}\neq z^{-1}_{j}$. (d) Sample $b\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$ and $o\leftarrow\mathsf{Dist}^{\mathsf{GenView}^{C},f}(j,z^{b},t)$ where $\mathsf{GenView}^{C}$ is Section 6.2 with oracle access to $C$, and $f(s,t=(\psi,t^{\prime}))=\mathsf{B}_{\psi}(s,t^{\prime})+\langle x_{\bar{{\cal H}}},\widetilde{y}_{\bar{{\cal H}}}\rangle$. # Recall that $x_{\bar{{\cal H}}}$ is a sub-string of $x_{-{\cal H}}$ (part of $t$) and that $\widetilde{y}_{\bar{{\cal H}}}$ is a sub-string of $\widetilde{y}_{-{\cal H}}$ (also part of $t$). (e) If $o=1$, output $b$. Otherwise, output $\bot$.
Proof of Property 2 of Section 5.1 using Section 6.2.
In the following, let ${\cal D}=\mathsf{GenView}^{C}()$ and $((X,U),(Y,V))\leftarrow C$. Consider a random execution of ${\tilde{\mathsf{B}}}^{\mathsf{B},C}(I,\>X_{-I},\>Y,\>V)$ for $I\leftarrow[n]$, and let $T,H,B,O,Z,J$ be the values of $\>\>t,h,{\cal H},o,z,j$ in the execution. Let $p={\mathrm{Pr}}_{(z,t)\leftarrow{\cal D},\>j\leftarrow[m]}\mathopen{}\mathclose{{}\left[\mathsf{Dist}^{{\cal D},f}(j,z^{(j)},t)=1}\right]$, and note that conditioned on $I\in H$, $J$ is distributed uniformly over $[m]$. Therefore,
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{Dist}^{{\cal D},f}(J,Z^{(J)},T)=1\mid I\in H}\right]=p$
and
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{Dist}^{{\cal D},f}(J,Z,T)=1\mid I\in H}\right]={\mathrm{Pr}}_{(z,t)\leftarrow{\cal D},\>j\leftarrow[m]}\mathopen{}\mathclose{{}\left[\mathsf{Dist}^{{\cal D},f}(j,z,t)=1}\right]\geq e^{\varepsilon}p+\frac{1}{m},$
where the inequality holds by Equation 29. Thus, the following holds for $X^{*}_{i}={\tilde{\mathsf{B}}}^{\mathsf{B},C}(i,\>X_{-i},\>Y,\>V)$:
${\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X^{*}_{i}=-X_{i}}\right]}\right]$ $={\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X_{I}^{*}=-X_{I}}\right]$
$={\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\{I\in H}\right\}\land\mathopen{}\mathclose{{}\left\{B=-X_{I}}\right\}\land\mathopen{}\mathclose{{}\left\{O=1}\right\}}\right]$
$=\mathopen{}\mathclose{{}\left(\frac{m}{n}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[H\neq\emptyset}\right]}\right)\cdot\frac{1}{2}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[O=1\mid\mathopen{}\mathclose{{}\left\{I\in H}\right\}\land\mathopen{}\mathclose{{}\left\{B=-X_{I}}\right\}}\right]$
$=\frac{m}{2n}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[H\neq\emptyset}\right]\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{Dist}^{{\cal D},f}(J,Z^{(J)},T)=1\mid I\in H}\right]$
$=\frac{m}{2n}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[H\neq\emptyset}\right]\cdot p,$
and
${\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X^{*}_{i}=X_{i}}\right]}\right]$ $={\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\{I\in H}\right\}\land\mathopen{}\mathclose{{}\left\{B=X_{I}}\right\}\land\mathopen{}\mathclose{{}\left\{O=1}\right\}}\right]$
$=\mathopen{}\mathclose{{}\left(\frac{m}{n}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[H\neq\emptyset}\right]}\right)\cdot\frac{1}{2}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[O=1\mid\mathopen{}\mathclose{{}\left\{I\in H}\right\}\land\mathopen{}\mathclose{{}\left\{B=X_{I}}\right\}}\right]$
$=\frac{m}{2n}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[H\neq\emptyset}\right]\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{Dist}^{{\cal D},f}(J,Z,T)=1\mid I\in H}\right]$
$>\frac{m}{2n}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[H\neq\emptyset}\right]\cdot(e^{\varepsilon}p+1/m)$
$=e^{\varepsilon}\cdot\mathopen{}\mathclose{{}\left(\frac{m}{2n}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[H\neq\emptyset}\right]\cdot p}\right)+\frac{1}{2n}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[H\neq\emptyset}\right],$
$>e^{\varepsilon}\cdot{\mathrm{E}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X^{*}_{i}=-X_{i}}\right]}\right]+\delta,$
which concludes the proof. The last inequality holds since ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[H\neq\emptyset}\right]\geq 1-\frac{1}{10000}$ and $\delta\leq\frac{1}{3n}$.
$\Box$
[…]
Appendix A [HaitnerMST22]’s Protocol Cannot Imply ${\sf OT}$
In this section, we show that for some carefully chosen ${\sf CDP}$ and accurate protocol $\Pi$, the joint view of the parties in [HaitnerMST22]’s protocol (Appendix A) can be simulated using a trivial protocol, without using $\Pi$ at all.
[…]
Oracle: An accurate ${\sf CDP}$ protocol $\Pi$ for the inner-product. Operation: 1. $\mathsf{A}$ and $\mathsf{B}$ choose random inputs $x\in\mathopen{}\mathclose{{}\left\{-1,1}\right\}^{n}$ and $y\in\mathopen{}\mathclose{{}\left\{-1,1}\right\}^{n}$, respectively. 2. The parties interact using $\Pi$ to get approximation $z$ of $\langle x,y\rangle$. 3. $\mathsf{A}$ chooses a random string $r\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}$, and sends $r,x_{r}=\mathopen{}\mathclose{{}\left\{x_{i}\colon r_{i}=1}\right\}$ to $\mathsf{B}$. $\mathsf{B}$ replies with $y_{-r}=\mathopen{}\mathclose{{}\left\{y_{i}\colon r_{i}=0}\right\}$. 4. Finally, $\mathsf{A}$ computes and outputs ${\rm out}_{\mathsf{A}}=\langle x_{-r},y_{-r}\rangle$, and $\mathsf{B}$ outputs ${\rm out}_{\mathsf{B}}=z-\langle x_{r},y_{r}\rangle$.
To see this, assume that $\Pi$ is a protocol that on inputs $x$ and $y$, outputs $z=\langle x,y\rangle+e_{\mathsf{A}}+e_{\mathsf{B}}$, where $e_{\mathsf{A}}$ and $e_{\mathsf{B}}$ are independent samples from the ${\rm Lap}(2/\varepsilon)$ distribution. Moreover, assume that $\Pi$ reveals $e_{\mathsf{A}}$ to $\mathsf{A}$ and $e_{\mathsf{B}}$ to $\mathsf{B}$ (and nothing else). Such a protocol is indeed differential private, and it can be implemented using secure multi-party computation. Moreover, as we show in this work, such a protocol can be used to construct OT. However, when composed with the ${\sf KA}$ protocol of [HaitnerMST22], the resulting protocol can be simulated trivially. [^14] More formally, and using the definition given in Section 4.4, we claim that the channel induced by executing Appendix A with oracle access to the channel $\Pi=\mathopen{}\mathclose{{}\left\{((x,(\langle x,y\rangle+e_{A}+e_{B},e_{A})),(y,(\langle x,y\rangle+e_{A}+e_{B},e_{B})))}\right\}_{x,y\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n},e_{A},e_{B}\leftarrow{\rm Lap}(2/\epsilon)}$ is a trivial channel.
Indeed, note that the view of $\mathsf{A}$ in $\Pi$ only contains $x,z$ and $e_{A}$, while the view of $\mathsf{B}$ only contains $y,z$ and $e_{B}$. In this case, the view of $\mathsf{A}$ in the ${\sf KA}$ protocol of [HaitnerMST22] contains $x,z,e_{A},r$ and $y_{-r}$, while the view of $\mathsf{B}$ contains $y,z,e_{B},r$ and $y_{r}$. We next explain how to simulate this view without using $\Pi$. Consider the following protocol $\Pi^{\prime}$ that simulates the ${\sf KA}$ protocol in a reverse order:
[…]
Operation: 1. $\mathsf{A}$ and $\mathsf{B}$ choose random inputs $x\in\mathopen{}\mathclose{{}\left\{-1,1}\right\}^{n}$ and $y\in\mathopen{}\mathclose{{}\left\{-1,1}\right\}^{n}$, respectively. 2. $\mathsf{A}$ chooses a random string $r\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}$, and sends $r,x_{r}=\mathopen{}\mathclose{{}\left\{x_{i}\colon r_{i}=1}\right\}$ to $\mathsf{B}$. $\mathsf{B}$ replies with $y_{-r}=\mathopen{}\mathclose{{}\left\{x_{i}\colon r_{i}=0}\right\}$. 3. $\mathsf{A}$ samples $e_{A}\leftarrow{\rm Lap}(2/\varepsilon)$ and sends $z_{A}=\langle x_{r},y_{r}\rangle+e_{A}$ to $\mathsf{B}$. $\mathsf{B}$ samples $e_{B}\leftarrow{\rm Lap}(2/\varepsilon)$ and sends $z_{B}=\langle x_{-r},y_{-r}\rangle+e_{B}$ to $\mathsf{A}$. 4. The output of the protocol is $z=z_{A}+z_{B}$.
Clearly, Appendix A is a trivial protocol and does not use any cryptographic assumptions. However, the views of $\mathsf{A}$ and $\mathsf{B}$ in $\Pi^{\prime}$ contain all the information learned by the parties in the ${\sf KA}$ protocol ($(x,z,e_{A},r,y_{-r})$ and $(y,z,e_{B},r,x_{r})$ respectively). Moreover, we claim that the parties in $\Pi^{\prime}$ do not learn any information that the parties in the ${\sf KA}$ protocol did not learn. Indeed, the only new value learned by $\mathsf{B}$, $z_{A},$ can be also computed by $\mathsf{B}$ in the ${\sf KA}$ protocol by computing $z-e_{B}-\langle x_{r},y_{r}\rangle$. Similarly, $z_{B}$ can be already computed by $\mathsf{A}$.
We note that every protocol that uses only a communication channel cannot be used to construct ${\sf OT}$ unless ${\sf OT}$ already exists, and similarly, every protocol that uses $\Pi^{\prime}$ as a subroutine (in a black-box manner) cannot be used to construct ${\sf OT}$. Since Appendix A (that is, the views and outputs of the parties when running Appendix A) are the same as $\Pi^{\prime}$, we conclude that Appendix A could not be used to construct ${\sf OT}$ in a black-box manner.
Appendix B WEC from AWEC (Proof of Section 4.4.4)
In this section, we prove Section 4.4.4 and show how to implement WEC (Section 4.4.3) from AWEC (Section 4.4.4) using a ppt protocol. Crucially, the security proof is constructive, so that it could be used in the computational case as well (see, Section 4.4.4).
To prove Section 4.4.4, we will need the following easy version of Goldriech-Levin [GoldreichL89].
[…]
There exists a ppt oracle-aided algorithm $\mathsf{Dec}$ such that the following holds. Let $n\in N$ be a number, $x\in{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}$, and and let $\mathsf{Pred}$ be an algorithm such that
${\mathrm{Pr}}_{r\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}}\mathopen{}\mathclose{{}\left[\mathsf{Pred}(r)=\operatorname{GL}(x,r)}\right]>3/4+0.001,$
where $\operatorname{GL}(x,r):=\langle x,r\rangle$ is the Goldreich-Levin predicate. Then ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{Dec}^{\mathsf{Pred}}(1^{n})=x}\right]=1-\operatorname{neg}(n)$.
Proof of Appendix B.
We use $\mathsf{Pred}$ to decode each bit of $x$ separately. For every $i$, let $e_{i}$ be the vector that has $1$ in the $i$-th entry, and 00’s everywhere else. Observe that, for a uniformly chosen $R\leftarrow{\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}}$,
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{Pred}(R)=\operatorname{GL}(x,R)\land\mathsf{Pred}(R\oplus e_{i})=\operatorname{GL}(x,R\oplus e_{i})}\right]\geq 1/2+0.001.$
Thus,
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{Pred}(R)\oplus\mathsf{Pred}(R\oplus e_{i})=\operatorname{GL}(x,R)\oplus\operatorname{GL}(x,R\oplus e_{i})}\right]\geq 1/2+0.001.$
By linearity of the inner product we get that,
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{Pred}(R)\oplus\mathsf{Pred}(R\oplus e_{i})=x_{i}}\right]\geq 1/2+0.001.$
Let $\mathsf{Dec}$ be the algorithm that for every $i$, computes $\mathsf{Pred}(R)\oplus\mathsf{Pred}(R\oplus e_{i})$ for $n$ random values of $R$, and let $x^{\prime}_{i}$ to be the majority of the outputs. Then, $\mathsf{Dec}$ outputs $x^{\prime}=x^{\prime}_{1},\dots,x^{\prime}_{n}$. By Chernoff bound, $x^{\prime}_{i}$ is equal to $x_{i}$ with all but negligible probability. By the union bound, the above is true for all $i$’s simultaneously with all but negligible probability, as we wanted to show.     $\Box$
We are now ready to prove Section 4.4.4.
Proof of Section 4.4.4.
We now define the protocol $\Lambda^{C}$ as follows:
[…]
Oracle access: A channel $C=((O_{\mathsf{A}},{V_{\mathsf{A}}}),(O_{\mathsf{B}},{V_{\mathsf{B}}}))$. Operation: 1. Sample $(({o_{\mathsf{A}}},{v_{\mathsf{A}}}),({o_{\mathsf{B}}},{v_{\mathsf{B}}}))\leftarrow C$. $\mathsf{A}$ gets $({o_{\mathsf{A}}},{v_{\mathsf{A}}})$ and $\mathsf{B}$ gets $({o_{\mathsf{B}}},{v_{\mathsf{B}}})$. 2. $\mathsf{A}$ chooses a random offset $s\leftarrow[1000\ell]$ and sends it to $\mathsf{B}$. Let ${o_{\mathsf{A}}}^{\prime}=\mathopen{}\mathclose{{}\left\lceil\frac{{o_{\mathsf{A}}}+s}{1000\ell}}\right\rceil$ and ${o_{\mathsf{B}}}^{\prime}=\mathopen{}\mathclose{{}\left\lceil\frac{{o_{\mathsf{B}}}+s}{1000\ell}}\right\rceil$ (if ${o_{\mathsf{B}}}=\bot$, let ${o_{\mathsf{B}}}^{\prime}=\bot$). 3. $\mathsf{A}$ chooses a random vector $r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{\log(n)}$ and sends it to $\mathsf{B}$. Let ${\widehat{o}_{\mathsf{A}}}=\langle{o_{\mathsf{A}}}^{\prime},r\rangle$ and ${\widehat{o}_{\mathsf{B}}}=\langle{o_{\mathsf{B}}}^{\prime},r\rangle$ (if ${o_{\mathsf{B}}}^{\prime}=\bot$, let ${\widehat{o}_{\mathsf{B}}}=\bot$). 4. $\mathsf{A}$ outputs ${\widehat{o}_{\mathsf{A}}}$ and $\mathsf{B}$ outputs ${\widehat{o}_{\mathsf{B}}}$.
Let $\tilde{C}$ be the channel induces by $\Lambda^{C}=(\mathsf{A},\mathsf{B})$ defined above. Let $O_{\mathsf{A}},{V_{\mathsf{A}}},O_{\mathsf{B}},{V_{\mathsf{B}}},S,R,O_{\mathsf{A}}^{\prime},O_{\mathsf{B}}^{\prime},{\widehat{O}_{\mathsf{A}}},{\widehat{O}_{\mathsf{B}}}$ be the random variables that corresponds to the value of ${o_{\mathsf{A}}},{v_{\mathsf{A}}},{o_{\mathsf{B}}},{v_{\mathsf{B}}},s,r,{o_{\mathsf{A}}}^{\prime},{o_{\mathsf{B}}}^{\prime},{\widehat{o}_{\mathsf{A}}},{\widehat{o}_{\mathsf{B}}}$ in a random execution of the above protocol. Denote ${\widehat{V}_{\mathsf{A}}}=({V_{\mathsf{A}}},S,R)$ and ${\widehat{V}_{\mathsf{B}}}=({V_{\mathsf{B}}},S,R)$ and note that $({\widehat{V}_{\mathsf{A}}},{\widehat{O}_{\mathsf{A}}}),({\widehat{V}_{\mathsf{B}}},{\widehat{O}_{\mathsf{B}}})$ defines the channels $\tilde{C}$.
We now prove that if $C=((O_{\mathsf{A}},{V_{\mathsf{A}}}),(O_{\mathsf{B}},{V_{\mathsf{B}}}))$ is an $(\ell,\alpha,p,q)$-AWEC then $\tilde{C}$ is an $(\alpha^{\prime}=\alpha+0.001,\>p^{\prime}=p,\>q^{\prime}=1/2+2.001q)$-WEC.
B.0.0.1 Agreement:
If $\mathopen{}\mathclose{{}\left|O_{\mathsf{A}}-O_{\mathsf{B}}}\right|\leq\ell$, then ${\mathrm{Pr}}_{S}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\lceil\frac{O_{\mathsf{A}}+S}{1000\ell}}\right\rceil\neq\mathopen{}\mathclose{{}\left\lceil\frac{O_{\mathsf{B}}+s}{1000\ell}}\right\rceil}\right]\leq 1/1000$. Thus,
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[{\widehat{O}_{\mathsf{A}}}\neq{\widehat{O}_{\mathsf{B}}}\mid{\widehat{O}_{\mathsf{B}}}\neq\bot}\right]$ $\leq{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[|O_{\mathsf{A}}-O_{\mathsf{B}}|\leq\ell\mid O_{\mathsf{B}}\neq\bot}\right]+1/1000$
$=\varepsilon+1/1000=\varepsilon^{\prime}$
B.0.0.2 $\mathsf{B}$’s privacy:
Recall that the view of $\mathsf{A}$ in the above protocol is ${\widehat{V}_{\mathsf{A}}}=({V_{\mathsf{A}}},S,R)$, and note that $S,R$ are independent of $O_{\mathsf{B}}$. Since, ${\widehat{O}_{\mathsf{B}}}=\bot$ iff $O_{\mathsf{B}}=\bot$ it follows that it follows that for every algorithm $\mathsf{D}$:
$\mathopen{}\mathclose{{}\left|{{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{D}({\widehat{V}_{\mathsf{A}}})=1\mid{\widehat{O}_{\mathsf{B}}}\neq\bot}\right]-{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{D}({\widehat{V}_{\mathsf{A}}})=1\mid{\widehat{O}_{\mathsf{B}}}=\bot}\right]}}\right|$
$=\mathopen{}\mathclose{{}\left|{{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{D}({V_{\mathsf{A}}},S,R)=1\mid O_{\mathsf{B}}\neq\bot}\right]-{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{D}({V_{\mathsf{A}}},S,R)=1\mid O_{\mathsf{B}}=\bot}\right]}}\right|$
$=\mathopen{}\mathclose{{}\left|{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{D}({V_{\mathsf{A}}})=1\mid O_{\mathsf{B}}\neq\bot}\right]-{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{D}({V_{\mathsf{A}}})=1\mid O_{\mathsf{B}}=\bot}\right]}\right|\leq p=p^{\prime}.$
B.0.0.3 $\mathsf{A}$’s privacy:
Assume towards a contradiction that there exists an algorithm $\mathsf{D}$ such that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{D}({\widehat{V}_{\mathsf{B}}})={\widehat{O}_{\mathsf{A}}}\mid{\widehat{O}_{\mathsf{B}}}\neq\bot}\right]\geq\frac{1+q^{\prime}}{2}=3/4+q+0.01.$ (30)
Let ${\cal G}=\mathopen{}\mathclose{{}\left\{({v_{\mathsf{B}}},s)\colon{\mathrm{Pr}}_{R}\mathopen{}\mathclose{{}\left[\mathsf{D}({v_{\mathsf{B}}},s,R)={\widehat{O}_{\mathsf{A}}}\mid{V_{\mathsf{B}}}={v_{\mathsf{B}}},S=s,{\widehat{O}_{\mathsf{B}}}\neq\bot}\right]\geq 3/4+0.001}\right\}$, and first note that ${\mathrm{Pr}}_{{V_{\mathsf{B}}},S}\mathopen{}\mathclose{{}\left[({V_{\mathsf{B}}},S)\in{\cal G}\mid{\widehat{O}_{\mathsf{B}}}\neq\bot}\right]\geq q+0.009$. Indeed, otherwise it holds that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{D}({\widehat{V}_{\mathsf{B}}})={\widehat{O}_{\mathsf{A}}}\mid{\widehat{O}_{\mathsf{B}}}\neq\bot}\right]$
$={\mathrm{Pr}}_{{V_{\mathsf{B}}},S}\mathopen{}\mathclose{{}\left[({V_{\mathsf{B}}},S)\in{\cal G}\mid{\widehat{O}_{\mathsf{B}}}\neq\bot}\right]+{\mathrm{Pr}}_{{V_{\mathsf{B}}},S}\mathopen{}\mathclose{{}\left[({V_{\mathsf{B}}},S)\notin{\cal G}\mid{\widehat{O}_{\mathsf{B}}}\neq\bot}\right]\cdot(3/4+0.001)$
$<(q+0.009)+1\cdot(3/4+0.001)$
$=3/4+q+0.01$
in contradiction to Equation 30. Next, by Appendix B there exists an algorithm $\mathsf{D}^{\prime}$ such that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{D}^{\prime}({V_{\mathsf{B}}},S)=O^{\prime}_{A}\mid({V_{\mathsf{B}}},S)\in{\cal G},O_{\mathsf{B}}^{\prime}\neq\bot}\right]\geq 1-o(1)$
Which implies that,
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{D}^{\prime}({V_{\mathsf{B}}},S)=O^{\prime}_{A}\mid O_{\mathsf{B}}^{\prime}\neq\bot}\right]$ $\geq{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{D}^{\prime}({V_{\mathsf{B}}},S)=O^{\prime}_{A}\mid({V_{\mathsf{B}}},S)\in{\cal G},O_{\mathsf{B}}^{\prime}\neq\bot}\right]\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{D}^{\prime}({V_{\mathsf{B}}},S)\in{\cal G}\mid{\widehat{O}_{\mathsf{B}}}\neq\bot}\right]$
$\geq(q+0.009)(1-o(1))$
$\geq q.$
Since by definition $\mathopen{}\mathclose{{}\left|(O^{\prime}_{A}\cdot 1000\ell-S)-O_{\mathsf{A}}}\right|\leq 1000\ell$, and $S$ is independent of ${V_{\mathsf{B}}}$, it follows that there exists an algorithm $\mathsf{D}^{\prime\prime}$ such that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|\mathsf{D}^{\prime\prime}({V_{\mathsf{B}}})-O_{\mathsf{A}}}\right|\leq 1000\ell\mid O_{\mathsf{B}}\neq\bot}\right]\geq\delta(1-2\alpha)>q.$
Contradicting the fact that $C$ is an $(\ell,\alpha,p,q)$-AWEC.     $\Box$
Appendix C Missing Proofs
C.1 Proving Sections 4.3 and 4.3
We make use of the following claim.
[…]
Let $X\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$ and let $X^{*}$ be a random variable over $\mathopen{}\mathclose{{}\left\{-1,1,\bot}\right\}$ (correlated with $X$) such that for every $b,b^{\prime}\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$:
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X^{*}=b\mid X=b^{\prime}}\right]\leq e^{\varepsilon}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X^{*}=b\mid X=-b^{\prime}}\right]+\delta.$
Then
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X^{*}=X}\right]\leq e^{\varepsilon}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X^{*}=-X}\right]+\delta.$
Proof.
Compute
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X^{*}=X}\right]$ $=\frac{1}{2}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X^{*}=-1\mid X=-1}\right]+\frac{1}{2}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X^{*}=1\mid X=1}\right]$
$\leq\frac{1}{2}\cdot\mathopen{}\mathclose{{}\left(e^{\varepsilon}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X^{*}=-1\mid X=1}\right]+\delta}\right)+\frac{1}{2}\cdot\mathopen{}\mathclose{{}\left(e^{\varepsilon}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X^{*}=1\mid X=-1}\right]+\delta}\right)$
$=e^{\varepsilon}\cdot\mathopen{}\mathclose{{}\left(\frac{1}{2}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X^{*}=-1\mid X=1}\right]+\frac{1}{2}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X^{*}=1\mid X=-1}\right]}\right)+\delta$
$=e^{\varepsilon}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X^{*}_{i}=-X_{i}}\right]+\delta.$
$\Box$
We next prove Section 4.3, restated below.
[…]
Let $f\colon\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}\rightarrow{\cal Y}$ be an $(\varepsilon,\delta)$-DP function, let $g\colon[n]\times\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n-1}\times{\cal Y}\rightarrow\mathopen{}\mathclose{{}\left\{-1,1,\bot}\right\}$ be a (randomized) function, and let $X=(X_{1},\ldots,X_{n})\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$. Then the following holds for every $i\in[n]$ where $X_{i}^{*}=g(i,X_{-i},f(X_{1},\ldots,X_{n}))$:
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X_{i}^{*}=X_{i}}\right]\leq e^{\varepsilon}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X_{i}^{*}=-X_{i}}\right]+\delta.$
Proof.
Fix $b,b^{\prime}\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$ and $i\in[n]$. By Section C.1, it is sufficient to prove that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X_{i}^{*}=b\mid X_{i}=b^{\prime}}\right]\leq e^{\varepsilon}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X_{i}^{*}=b\mid X_{i}=-b^{\prime}}\right]+\delta.$ (31)
For $x_{-i}=(x_{1},\ldots,x_{i-1},x_{i+1},\ldots,x_{n})\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n-1}$, define the function
$h_{x_{-i}}(y)=g(i,x_{-i},y).$
Since $f$ is $(\varepsilon,\delta)$-DP, for any $x_{-i}\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n-1}$ it holds that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[h_{x_{-i}}(f(x_{1},\ldots,x_{i-1},b^{\prime},x_{i+1},\ldots,x_{n}))=b}\right]\leq e^{\varepsilon}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[h_{x_{-i}}(f(x_{1},\ldots,x_{i-1},-b^{\prime},x_{i+1},\ldots,x_{n}))=b}\right]+\delta.$
Thus,
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X_{i}^{*}=b\mid X_{i}=b^{\prime}}\right]$ $={\mathrm{E}}_{x_{-i}\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n-1}}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[g(i,x_{-i},f(x_{1},\ldots,x_{i-1},b^{\prime},x_{i+1},\ldots,x_{n}))=b}\right]}\right]$
$={\mathrm{E}}_{x_{-i}\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n-1}}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[h_{x_{-i}}(f(x_{1},\ldots,x_{i-1},b^{\prime},x_{i+1},\ldots,x_{n}))=b}\right]}\right]$
$\leq e^{\varepsilon}\cdot{\mathrm{E}}_{x_{-i}\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n-1}}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[h_{x_{-i}}(f(x_{1},\ldots,x_{i-1},-b^{\prime},x_{i+1},\ldots,x_{n}))=b}\right]}\right]+\delta$
$=e^{\varepsilon}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X_{i}^{*}=b\mid X_{i}=-b^{\prime}}\right]+\delta.$
$\Box$
We next prove Section 4.3, restated below.
[…]
Let $f=\mathopen{}\mathclose{{}\left\{f_{\kappa}\colon\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n(\kappa)}\rightarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{m(\kappa)}}\right\}_{\kappa\in{\mathbb{N}}}$ be an $(\varepsilon,\delta)$-CDP function ensemble, and let $\mathopen{}\mathclose{{}\left\{g_{\kappa}}\right\}_{\kappa\in{\mathbb{N}}}$ be a poly-size circuit family. Then, for large enough $\kappa$ and $X=(X_{1},\ldots,X_{n(\kappa)})\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n(\kappa)}$, the following holds for every $i\in[n(\kappa)]$ where $X_{i}^{*}=g_{\kappa}(i,X_{-i},f_{\kappa}(X_{1},\ldots,X_{n}))$:
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X_{i}^{*}=X_{i}}\right]\leq e^{\varepsilon(\kappa)}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X_{i}^{*}=-X_{i}}\right]+\delta(\kappa).$
Proof.
In the following, fix $b,b^{\prime}\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$, and for $\kappa\in{\mathbb{N}}$, $i\in[n(\kappa)]$ and $x_{-i}\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n(\kappa)-1}$, define
$h_{\kappa}^{i,x_{-i}}(y)=b\cdot g_{\kappa}(i,x_{-i},y).$
Note that for any ensemble $\mathopen{}\mathclose{{}\left\{(i,x_{-i})_{\kappa}\in[n(\kappa)]\times\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n(\kappa)-1}}\right\}_{\kappa\in{\mathbb{N}}}$, the circuit family
$\mathopen{}\mathclose{{}\left\{h_{\kappa}^{(i,x_{-i})_{\kappa}}=g_{\kappa}((i,x_{-i})_{\kappa},\cdot)}\right\}_{\kappa\in{\mathbb{N}}}$ has poly-size. Since $f=\mathopen{}\mathclose{{}\left\{f_{\kappa}}\right\}_{\kappa\in{\mathbb{N}}}$ is $(\varepsilon,\delta)$-CDP, then for large enough $\kappa$, the following holds for every $i\in[n(\kappa)]$ and $x_{-i}\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n(\kappa)-1}$:
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[h_{\kappa}^{i,x_{-i}}(f_{\kappa}(x_{1},\ldots,x_{i-1},b^{\prime},x_{i+1},\ldots,x_{n}))=1}\right]$
$\leq e^{\varepsilon(\kappa)}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[h_{\kappa}^{i,x_{-i}}(f(x_{1},\ldots,x_{i-1},-b^{\prime},x_{i+1},\ldots,x_{n}))=1}\right]+\delta(\kappa),$
as otherwise, there would exist an ensemble $\mathopen{}\mathclose{{}\left\{(i,x_{-i})_{\kappa}\in[n(\kappa)]\times\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n(\kappa)-1}}\right\}_{\kappa\in{\cal S}}$ for an infinite set ${\cal S}\subseteq{\mathbb{N}}$ such that the circuit family $\mathopen{}\mathclose{{}\left\{h_{\kappa}^{(i,x_{-i})_{\kappa}}}\right\}_{\kappa\in{\cal S}}$ violates the $(\varepsilon,\delta)$-CDP property of $f$.
In the following, fix such large enough $\kappa$ and $i\in[n]$ for $n=n(\kappa)$, let $X=(X_{1},\ldots,X_{n})\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}$ and $X_{i}^{*}=g_{\kappa}(i,X_{-i},f_{\kappa}(X_{1},\ldots,X_{n}))$, and compute
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X_{i}^{*}=b\mid X_{i}=b^{\prime}}\right]$
$={\mathrm{E}}_{x_{-i}\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n-1}}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[g_{\kappa}(i,x_{-i},f_{\kappa}(x_{1},\ldots,x_{i-1},b^{\prime},x_{i+1},\ldots,x_{n}))=b}\right]}\right]$
$={\mathrm{E}}_{x_{-i}\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n-1}}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[h_{\kappa}^{i,x_{-i}}(f(x_{1},\ldots,x_{i-1},b^{\prime},x_{i+1},\ldots,x_{n}))=1}\right]}\right]$
$\leq e^{\varepsilon(\kappa)}\cdot{\mathrm{E}}_{x_{-i}\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n-1}}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[h_{\kappa}^{i,x_{-i}}(f(x_{1},\ldots,x_{i-1},-b^{\prime},x_{i+1},\ldots,x_{n}))=1}\right]}\right]+\delta(\kappa)$
$=e^{\varepsilon(\kappa)}\cdot{\mathrm{E}}_{x_{-i}\leftarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n-1}}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[g_{\kappa}(i,x_{-i},f_{\kappa}(x_{1},\ldots,x_{i-1},-b^{\prime},x_{i+1},\ldots,x_{n}))=b}\right]}\right]+\delta(\kappa)$
$=e^{\varepsilon(\kappa)}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X_{i}^{*}=b\mid X_{i}=-b^{\prime}}\right]+\delta(\kappa).$
Since the above holds for any $b,b^{\prime}\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$, we conclude by Section C.1 that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X_{i}^{*}=X_{i}}\right]\leq e^{\varepsilon(\kappa)}\cdot{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[X_{i}^{*}=-X_{i}}\right]+\delta(\kappa),$
as required.     $\Box$
C.2 Proving Section 6.1
To prove Section 6.1, we use the following lemma that measures the distance between two uniform stings conditioned on a random index $i$ either being fixed to 00 or to $1$.
[…]
Let $R\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}$. For any (randomized) function $F:\{0,1\}^{n}\rightarrow\{0,1\}$ and $\alpha>0$, it holds that
${\mathrm{Pr}}_{i\leftarrow[n]}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|\>{\mathrm{E}}\mathopen{}\mathclose{{}\left[F(R)\mid R_{i}=0}\right]-{\mathrm{E}}\mathopen{}\mathclose{{}\left[F(R)\mid R_{i}=1}\right]\>}\right|\geq\alpha}\right]\leq\frac{2}{n\alpha^{2}},$ (32)
where the expectations are taken over $R$ and the randomness of $F$.
The proof of Section C.2 uses the following fact.
[…]
Let $R$ be uniform random variable over $\{0,1\}^{n}$, and let $I$ be uniform random variable over $\mathcal{I}\subseteq[n]$, independent of $R$, then $SD(R|_{R_{I}=0},R|_{R_{I}=1})\leq 1/\sqrt{\mathopen{}\mathclose{{}\left|{\cal I}}\right|}$.
We first prove Section C.2 using Section C.2.
Proof of Section C.2.
Assume towards a contradiction that there exist $F$ and $\alpha$ such that Equation 32 does not hold. Namely, for $m=1/\alpha^{2}$, there exist more than $2m$ indices $i\in[n]$ with
$\mathopen{}\mathclose{{}\left|\>{\mathrm{E}}\mathopen{}\mathclose{{}\left[F(R)\mid R_{i}=0}\right]-{\mathrm{E}}\mathopen{}\mathclose{{}\left[F(R)\mid R_{i}=1}\right]\>}\right|\geq 1/\sqrt{m}.$
This implies that there exist $b\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$ and more than $m$ indices $i\in[n]$ with
${\mathrm{E}}\mathopen{}\mathclose{{}\left[F(R)\mid R_{i}=b}\right]-{\mathrm{E}}\mathopen{}\mathclose{{}\left[F(R)\mid R_{i}=1-b}\right]\geq 1/\sqrt{m}$
(denote this set by ${\cal I}$). Thus, we deduce for $I\leftarrow{\cal I}$ that
$\mathopen{}\mathclose{{}\left|{\mathrm{E}}\mathopen{}\mathclose{{}\left[F(R)\mid R_{I}=0}\right]-{\mathrm{E}}\mathopen{}\mathclose{{}\left[F(R)\mid R_{I}=1}\right]}\right|$ $\geq{\mathrm{E}}\mathopen{}\mathclose{{}\left[F(R)\mid R_{I}=b}\right]-{\mathrm{E}}\mathopen{}\mathclose{{}\left[F(R)\mid R_{I}=1-b}\right]$ (33)
$\geq 1/\sqrt{m}.$
On the other hand, note that
$SD(F(R)|_{R_{I}=0},F(R)|_{R_{I}=1})\leq SD(R|_{R_{I}=0},R|_{R_{I}=1})\leq 1/\sqrt{\mathopen{}\mathclose{{}\left|{\cal I}}\right|}<1/\sqrt{m},$
where the second inequality holds by Section C.2. Thus, we conclude that
$\mathopen{}\mathclose{{}\left|{\mathrm{E}}\mathopen{}\mathclose{{}\left[F(R)\mid R_{I}=0}\right]-{\mathrm{E}}\mathopen{}\mathclose{{}\left[F(R)\mid R_{I}=1}\right]}\right|$ $\leq SD(F(R)|_{R_{I}=0},F(R)|_{R_{I}=1})\cdot\sup_{r\in\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n},s\in\mathopen{}\mathclose{{}\left\{0,1}\right\}^{*}}(\mathopen{}\mathclose{{}\left|F_{s}(r)}\right|)$
$<1/\sqrt{m},$
where $F_{s}(r)$ denotes the function $F$ when fixing its random coins to $s$. The above inequality contradicts Equation 33, concluding the proof of the lemma.     $\Box$
Using Section C.2, we now prove Section 6.1, restated below.
[…]
Let $(Z,W)\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{0,1}\right\}^{*}$ be jointly distributed random variables, let $R\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}$ and $I\leftarrow[n]$ (sampled independently), and let $F$ be a (randomized) function that satisfies
$\mathopen{}\mathclose{{}\left|{\mathrm{E}}\mathopen{}\mathclose{{}\left[F(R,Z_{R},W)-F(R,Z^{(I)}_{R},W)}\right]}\right|\geq\gamma,$
for $Z^{(I)}=(Z_{1},\ldots,Z_{I-1},-Z_{I},Z_{I+1},\ldots,Z_{n})$. Then
  1. 1.
    ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}^{F}(I,Z_{-I},W)=-Z_{I}}\right]\leq O\mathopen{}\mathclose{{}\left(\frac{1}{\gamma^{2}n}}\right)$, and
  2. 2.
    ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}^{F}(I,Z_{-I},W)=Z_{I}}\right]\geq\Omega(\gamma)-O\mathopen{}\mathclose{{}\left(\frac{1}{\gamma^{2}n}}\right)$.
Proof.
Let $\gamma\in(0,1)$ and $n\in{\mathbb{N}}$ and consider the following algorithm $\mathsf{G}=\mathsf{G}_{\gamma}$:
[…]
Inputs: $i\in[n]$, $z_{-i}=(z_{1},\ldots,z_{i-1},z_{i+1},\ldots,z_{n})\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n-1}$ and $w\in\mathopen{}\mathclose{{}\left\{0,1}\right\}^{*}$. Parameter: $\gamma\in(0,1)$. Oracle: $F\colon\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{\leq n}\times\mathopen{}\mathclose{{}\left\{0,1}\right\}^{*}\rightarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}$. Operation: 1. For $b\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$: (a) Let $z^{b}=(z_{1},\ldots,z_{i-1},b,z_{i+1},\ldots,z_{n})$. (b) Estimate $\mu^{b}:={\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\mid r_{i}=1,\>F}\mathopen{}\mathclose{{}\left[F(r,z^{b}_{r},w)}\right]$ as follows: Sample $r_{1},\ldots,r_{s}\leftarrow\mathopen{}\mathclose{{}\left\{r\in\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\colon r_{i}=1}\right\}$, for $s=\mathopen{}\mathclose{{}\left\lceil\frac{128\log(12n)}{\gamma^{2}}}\right\rceil$, and then sample $\tilde{\mu}^{b}\leftarrow\frac{1}{s}\sum_{j=1}^{s}F(r_{j},z^{b}_{r_{j}},w)$ (using $s$ oracle calls to $F$). 2. Estimate $\mu^{*}:={\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\mid r_{i}=0,\>f}\mathopen{}\mathclose{{}\left[f(r,z^{1}_{r},w)}\right]$ as follows: Sample $r_{1},\ldots,r_{s}\leftarrow\mathopen{}\mathclose{{}\left\{r\in\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\colon r_{i}=0}\right\}$, for $s=\mathopen{}\mathclose{{}\left\lceil\frac{128\log(12n)}{\gamma^{2}}}\right\rceil$, and then sample $\tilde{\mu}^{*}\leftarrow\frac{1}{s}\sum_{j=1}^{s}F(r_{j},z^{1}_{r_{j}},w)$ (using $s$ oracle calls to $F$). 3. If exists $b\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$ s.t. $\mathopen{}\mathclose{{}\left|\tilde{\mu}^{b}-\tilde{\mu}^{*}}\right|<\gamma/4$ and $\mathopen{}\mathclose{{}\left|\tilde{\mu}^{-b}-\tilde{\mu}^{*}}\right|>\gamma/4$, output $b$. 4. Otherwise, output $\bot$.
In the following, fix a pair of (jointly distributed) random variables $(Z,W)\in\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{0,1}\right\}^{*}$ and a randomized function $F\colon\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\times\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}^{\leq n}\times\mathopen{}\mathclose{{}\left\{0,1}\right\}^{*}\rightarrow\mathopen{}\mathclose{{}\left\{\!\shortminus 1,1\!}\right\}$ that satisfy
$\mathopen{}\mathclose{{}\left|{\mathrm{E}}\mathopen{}\mathclose{{}\left[F(R,Z_{R},W)-F(R,Z^{(I)}_{R},W)}\right]}\right|\geq\gamma,$
for $R\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}$ and $I\leftarrow[n]$ that are sampled independently. Our goal is to prove that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}^{F}(I,Z_{-I},W)=-Z_{I}}\right]\leq O\mathopen{}\mathclose{{}\left(\frac{1}{\gamma^{2}n}}\right),$ (34)
and
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}^{F}(I,Z_{-I},W)=Z_{I}}\right]\geq\Omega(\gamma)-O\mathopen{}\mathclose{{}\left(\frac{1}{\gamma^{2}n}}\right).$ (35)
Note that
$\text{For every random variable }D\in[-1,1]\text{ with }\mathopen{}\mathclose{{}\left|{\mathrm{E}}\mathopen{}\mathclose{{}\left[D}\right]}\right|\geq\gamma>0:\quad{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|D}\right|>\gamma/2}\right]>\gamma/2,$ (36)
as otherwise, $\mathopen{}\mathclose{{}\left|{\mathrm{E}}\mathopen{}\mathclose{{}\left[D}\right]}\right|\leq{\mathrm{E}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|D}\right|}\right]\leq 1\cdot\frac{\gamma}{2}+\frac{\gamma}{2}(1-\frac{\gamma}{2})<\gamma$.
By applying Equation 36 with $D={\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n},F}\mathopen{}\mathclose{{}\left[F(r,Z_{r},W)-f(r,Z^{(I)}_{r},W)}\right]$, it holds that
${\mathrm{Pr}}_{(i,z,w)\leftarrow(I,Z,W)}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|{\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n},F}\mathopen{}\mathclose{{}\left[F(r,z_{r},w)-F(r,z^{(i)}_{r},w)}\right]}\right|>\gamma/2}\right]>\gamma/2.$ (37)
On the other hand, for every fixing of $(z,w)\in\operatorname{Supp}(Z,W)$, we can apply Section C.2 with the function $F_{z,w}(r)=F(r,z_{r},w)$ and with $\alpha=\frac{\gamma}{16}$ to obtain that
${\mathrm{Pr}}_{i\leftarrow I}\mathopen{}\mathclose{{}\left[\>\mathopen{}\mathclose{{}\left|{\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\mid r_{i}=0,\>F}\mathopen{}\mathclose{{}\left[F(r,z_{r},w)}\right]-{\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\mid r_{i}=1,\>F}\mathopen{}\mathclose{{}\left[F(r,z_{r},w)}\right]\>}\right|\geq\frac{\gamma}{16}}\right]\leq\frac{512}{n\gamma^{2}}.$
But since the above holds for every fixing of $(z,w)$, then in particular it holds that
${\mathrm{Pr}}_{(i,z,w)\leftarrow(I,Z,W)}\mathopen{}\mathclose{{}\left[\>\mathopen{}\mathclose{{}\left|{\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\mid r_{i}=0,\>F}\mathopen{}\mathclose{{}\left[F(r,z_{r},w)}\right]-{\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\mid r_{i}=1,\>F}\mathopen{}\mathclose{{}\left[F(r,z_{r},w)}\right]\>}\right|\geq\frac{\gamma}{16}}\right]\leq\frac{512}{n\gamma^{2}}.$ (38)
We next prove Equations 34 and 35 using Equations 37 and 38.
In the following, for a triplet $t=(i,z,w)\in\operatorname{Supp}(I,Z,W)$, consider a random execution of $\mathsf{G}^{F}(i,z_{-i},w)$. For $x\in\mathopen{}\mathclose{{}\left\{-1,1,*}\right\}$, let $\mu^{x}_{t}$ be the value of $\mu^{x}$ in the execution, and let $M^{x}_{t}$ be the (random variable of the) value of $\tilde{\mu}^{x}$ in the execution. Note that by definition, it holds that $\mu_{i,z,w}^{z_{i}}={\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n},F}\mathopen{}\mathclose{{}\left[F(r,z_{r},w)}\right]$ and $\mu_{i,z,w}^{-z_{i}}={\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n},F}\mathopen{}\mathclose{{}\left[F(r,z^{(i)}_{r},w)}\right]$. Therefore, Equation 37 is equivalent to
${\mathrm{Pr}}_{t=(i,z,w)\leftarrow(I,Z,W)}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|\mu_{t}^{z_{i}}-\mu_{t}^{-z_{i}}}\right|>\gamma/2}\right]>\gamma/2.$ (39)
Furthermore, note that $\mu_{i,z,w}^{*}:={\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\mid r_{i}=0,\>F}\mathopen{}\mathclose{{}\left[F(r,z^{1}_{r},w)}\right]={\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\mid r_{i}=0,\>F}\mathopen{}\mathclose{{}\left[F(r,z_{r},w)}\right]$ and that $\mu_{i,z,w}^{z_{i}}={\mathrm{E}}_{r\leftarrow\mathopen{}\mathclose{{}\left\{0,1}\right\}^{n}\mid r_{i}=1,\>F}\mathopen{}\mathclose{{}\left[F(r,z^{z_{i}}_{r},w)}\right]$. Therefore, Equation 38 is equivalent to
${\mathrm{Pr}}_{t=(i,z,w)\leftarrow(I,Z,W)}\mathopen{}\mathclose{{}\left[\>\mathopen{}\mathclose{{}\left|\mu_{t}^{*}-\mu_{t}^{z_{i}}}\right|\geq\frac{\gamma}{16}}\right]\leq\frac{512}{n\gamma^{2}}.$ (40)
We next prove the lemma using Equations 39 and 40.
Note that by Hoeffding’s inequality, for every $t=(i,z,w)\in\operatorname{Supp}(I,Z,W)$ and $x\in\mathopen{}\mathclose{{}\left\{-1,1,*}\right\}$ it holds that ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|M_{t}^{x}-\mu_{t}^{x}}\right|\geq\frac{\gamma}{16}}\right]\leq 2\cdot e^{-2s\mathopen{}\mathclose{{}\left(\frac{\gamma}{16}}\right)^{2}}\leq\frac{1}{6n}$, which yields that for every fixing of $t=(i,z,w)\in\operatorname{Supp}(I,Z,W)$, w.p. at least $1-\frac{1}{2n}$ we have for all $x\in\mathopen{}\mathclose{{}\left\{-1,1,*}\right\}$ that $\mathopen{}\mathclose{{}\left|M_{t}^{x}-\mu_{t}^{x}}\right|<\frac{\gamma}{16}$ (denote this event by $E_{t}$).
The proof of Equation 34 holds by the following calculation:
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}^{F}(I,Z_{-I},W)=-Z_{I}}\right]$
$={\mathrm{E}}_{(i,z,w)\leftarrow(I,Z,W)}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}^{F}(i,z_{-i},w)=-z_{i}}\right]}\right]$
$={\mathrm{E}}_{t=(i,z,w)\leftarrow(I,Z,W)}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left|M_{t}^{*}-M_{t}^{z_{i}}}\right|>\gamma/4}\right\}\land\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left|M_{t}^{*}-M_{t}^{-z_{i}}}\right|<\gamma/4}\right\}}\right]}\right]$
$\leq{\mathrm{E}}_{t=(i,z,w)\leftarrow(I,Z,W)}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|M_{t}^{*}-M_{t}^{z_{i}}}\right|>\gamma/4}\right]}\right]$
$\leq{\mathrm{E}}_{t=(i,z,w)\leftarrow(I,Z,W)}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|M_{t}^{*}-M_{t}^{z_{i}}}\right|>\gamma/4\mid E_{t}}\right]}\right]+\frac{1}{2n}$
$\leq{\mathrm{Pr}}_{t=(i,z,w)\leftarrow(I,Z,W)}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left|\mu_{t}^{*}-\mu_{t}^{z_{i}}}\right|\geq\frac{\gamma}{16}}\right]+\frac{1}{2n}$
$\leq\frac{512}{n\gamma^{2}}+\frac{1}{2n},$
The second inequality holds since ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\neg E_{t}}\right]\leq\frac{1}{2n}$ for every $t$. The penultimate inequality holds since conditioned on $E_{t}$, it holds that $\mathopen{}\mathclose{{}\left|M_{t}^{*}-\mu_{t}^{*}}\right|\leq\frac{\gamma}{16}$ and $\mathopen{}\mathclose{{}\left|M_{t}^{z_{i}}-\mu_{t}^{z_{i}}}\right|\leq\frac{\gamma}{16}$, which implies that $\mathopen{}\mathclose{{}\left|M_{t}^{*}-M_{t}^{z_{i}}}\right|>\gamma/4\>\implies\>\mathopen{}\mathclose{{}\left|\mu_{t}^{*}-\mu_{t}^{z_{i}}}\right|\geq\frac{\gamma}{4}-2\cdot\frac{\gamma}{16}>\frac{\gamma}{16}$. The last inequality holds by Equation 40.
It is left to prove Equation 35. Observe that Equations 39 and 40 imply with probability at least $\gamma/2-\frac{512}{n\gamma^{2}}$ over $t=(i,z,w)\in(I,Z,W)$, we have $\mathopen{}\mathclose{{}\left|\mu_{t}^{*}-\mu_{t}^{z_{i}}}\right|\leq\frac{\gamma}{16}$ and $\mathopen{}\mathclose{{}\left|\mu_{t}^{z_{i}}-\mu_{t}^{-z_{i}}}\right|\geq\frac{\gamma}{2}$. Therefore, conditioned on event $E_{t}$ (i.e., $\forall x\in\mathopen{}\mathclose{{}\left\{-1,1,*}\right\}:\>\mathopen{}\mathclose{{}\left|M_{t}^{x}-\mu_{t}^{x}}\right|\leq\frac{\gamma}{16}$), we have for such triplets $t=(i,z,w)$ that
$\mathopen{}\mathclose{{}\left|M_{t}^{*}-M_{t}^{-z_{i}}}\right|$ $\geq\mathopen{}\mathclose{{}\left|\mu_{t}^{z_{i}}-\mu_{t}^{-z_{i}}}\right|-\mathopen{}\mathclose{{}\left|\mu_{t}^{z_{i}}-\mu_{t}^{*}}\right|-\mathopen{}\mathclose{{}\left|M_{t}^{*}-\mu_{t}^{*}}\right|-\mathopen{}\mathclose{{}\left|M_{t}^{-z_{i}}-\mu_{t}^{-z_{i}}}\right|$
$\geq\frac{\gamma}{2}-3\cdot\frac{\gamma}{16}$
$>\gamma/4,$
and
$\mathopen{}\mathclose{{}\left|M_{t}^{*}-M_{t}^{z_{i}}}\right|$ $\leq\mathopen{}\mathclose{{}\left|M_{t}^{*}-\mu_{t}^{*}}\right|+\mathopen{}\mathclose{{}\left|\mu_{t}^{*}-\mu_{t}^{z_{i}}}\right|+\mathopen{}\mathclose{{}\left|\mu_{t}^{z_{i}}-M_{t}^{z_{i}}}\right|$
$\leq 3\cdot\frac{\gamma}{16}$
$<\gamma/4.$
Thus, we conclude that
${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}^{F}(I,Z_{-I},W)=Z_{I}}\right]$
$={\mathrm{E}}_{(i,z,w)\leftarrow(I,Z,W)}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathsf{G}^{F}(i,z_{-i},w)=z_{i}}\right]}\right]$
$={\mathrm{E}}_{t=(i,z,w)\leftarrow(I,Z,W)}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left|M_{t}^{*}-M_{t}^{-z_{i}}}\right|>\gamma/4}\right\}\land\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left|M_{t}^{*}-M_{t}^{z_{i}}}\right|<\gamma/4}\right\}}\right]}\right]$
$\geq\mathopen{}\mathclose{{}\left(1-\frac{1}{2n}}\right)\cdot{\mathrm{E}}_{t=(i,z,w)\leftarrow(I,Z,W)}\mathopen{}\mathclose{{}\left[{\mathrm{Pr}}\mathopen{}\mathclose{{}\left[\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left|M_{t}^{*}-M_{t}^{-z_{i}}}\right|>\gamma/4}\right\}\land\mathopen{}\mathclose{{}\left\{\mathopen{}\mathclose{{}\left|M_{t}^{*}-M_{t}^{z_{i}}}\right|<\gamma/4}\right\}\mid E_{t}}\right]}\right]$
$\geq\mathopen{}\mathclose{{}\left(1-\frac{1}{2n}}\right)\cdot\mathopen{}\mathclose{{}\left(\gamma/2-\frac{512}{n\gamma^{2}}}\right)$
$\geq\gamma/4-\frac{512}{n\gamma^{2}},$
which proves Equation 35. The first inequality holds since ${\mathrm{Pr}}\mathopen{}\mathclose{{}\left[E_{t}}\right]\geq 1-\frac{1}{2n}$ for every $t=(i,z,w)$, and the second one holds by the observation above.
$\Box$